### v3

# Xing Huang(a), Eva Kopfer(b), Panpan Ren(c) a)

Abstract
The exponential ergodicity in the $L^{1}$-Wasserstein distance for partially dissipative McKean-Vlasov SDEs has been extensively studied. However, the question of exponential ergodicity in the $L^{2}$-Wasserstein distance and relative entropy has remained unresolved. This paper addresses the problem by establishing the log-Sobolev inequality for both the time-marginal distributions and the invariant probability measure, providing a positive resolution. As part of the groundwork, the log-Sobolev inequality is investigated for the associated time-inhomogeneous semigroup. The main results are further extended to degenerate diffusion.
AMS subject Classification:  60H10, 60K35, 82C22.
Keywords: Log-Sobolev inequality, Dimension-free Harnack inequality, Talagrand inequality, McKean-Vlasov SDEs, exponential ergodicity, $L^{2}$-Wasserstein distance
1 Introduction
We aim to advance the study of functional inequalities in McKean–Vlasov SDEs and their corresponding decoupled equations. The McKean–Vlasov SDE, first introduced by McKean [23], was developed to study the nonlinear Fokker-Planck equation. The idea goes back to Kac [18, 19], who characterized nonlinear PDEs in Vlasov’s kinetic theory. The main tool is the so-called “propagation of chaos” which builds a corner stone in mean field interacting particle systems, see e.g. [9, 24, 31]. The McKean–Vlasov SDE is the prototype of a distribution–dependent stochastic differential equation (DDSDE).
In general, the McKean–Vlasov SDE is given by
(1.1) $\text{\rm{d}}X_{t}=b_{t}(X_{t},\mathscr{L}_{X_{t}})\text{\rm{d}}t+\sigma_{t}(X_{t},\mathscr{L}_{X_{t}})\text{\rm{d}}W_{t},$
where $\{W_{t}\}_{t\geq 0}$ is an $n$-dimensional Brownian motion on some complete filtered probability space $(\Omega,\mathscr{F},(\mathscr{F}_{t})_{t\geq 0},\mathbb{P})$ and $\mathscr{L}_{X_{t}}$ is the distribution of $X_{t}$. The coefficients $b:[0,\infty)\times\mathbb{R}^{d}\times\mathscr{P}(\mathbb{R}^{d})\rightarrow\mathbb{R}^{d}$, $\sigma:[0,\infty)\times\mathbb{R}^{d}\times\mathscr{P}(\mathbb{R}^{d})\rightarrow\mathbb{R}^{d}\otimes\mathbb{R}^{n}$ are assumed to be measurable and bounded on bounded sets, where $\mathscr{P}(\mathbb{R}^{d})$ is the set of all probability measures on $\mathbb{R}^{d}$ equipped with the weak topology.
In general, we will consider the solution of (1.1) in a subspace of $\mathscr{P}(\mathbb{R}^{d})$. To this end, for $p\geq 1$, let
$\mathscr{P}_{p}(\mathbb{R}^{d}):=\big{\{}\mu\in\mathscr{P}(\mathbb{R}^{d}):\|\mu\|_{p}:=\mu(|\cdot|^{p})^{\frac{1}{p}}<\infty\big{\}},$
which is a Polish space under the $L^{p}$-Wasserstein distance
$\mathbb{W}_{p}(\mu,\nu)=\inf_{\pi\in\mathscr{C}(\mu,\nu)}\bigg{(}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}|x-y|^{p}\pi(\text{\rm{d}}x,\text{\rm{d}}y)\bigg{)}^{\frac{1}{p}},\ \ \mu,\nu\in\mathscr{P}_{p}(\mathbb{R}^{d}),$
where $\mathscr{C}(\mu,\nu)$ is the set of all couplings of $\mu$ and $\nu$.
The well-posedness for (1.1) in the weak and strong sense has been intensively investigated, see e.g. [36, 39] and references within. When (1.1) is well-posed in $\mathscr{P}_{p}(\mathbb{R}^{d})$, for given $\gamma\in\mathscr{P}_{p}(\mathbb{R}^{d})$ we denote $P_{t}^{*}\gamma=\mathscr{L}_{X_{t}}$ with $\mathscr{L}_{X_{0}}=\gamma$.
Recall that the relative entropy between $\mu,\nu\in\mathscr{P}(\mathbb{R}^{d})$ is defined as
$\mathrm{Ent}(\nu|\mu)=\left\{\begin{array}[]{ll}\nu(\log(\frac{\text{\rm{d}}\nu}{\text{\rm{d}}\mu})),&\hbox{$\nu$ \text{is absolutely continuous w.r.t.} $\mu$;}\\ \infty,&\hbox{otherwise.}\end{array}\right.$
In the case $b_{t}=b,\sigma_{t}=I_{d\times d}$, exponential ergodicity of a system guarantees the existence of an invariant probability measure with exponential rate of convergence measured in some metric. For the McKean-Vlasov SDE, the $L^{2}$-Wasserstein distance and the relative entropy are frequently used. Provided that the coefficient is uniformly dissipative, i.e.
$\langle b(x,\mu)-b(y,\tilde{\mu}),x-y\rangle\leq-\alpha|x-y|^{2}+\beta\mathbb{W}_{2}(\mu,\tilde{\mu})^{2},\ \ x,y\in\mathbb{R}^{d},\mu,\tilde{\mu}\in\mathscr{P}_{2}(\mathbb{R}^{d}),$
for some $\alpha>\beta>0$, exponential ergodicity holds in the $L^{2}$-Wasserstein distance as well as in relative entropy, see [17, 30, 36] and references therein. In the partially dissipative case, i.e. dissipation only occurs at long distances, exponential ergodicity has been established in the $L^{1}$-Wasserstein distance, see [11, 16, 21, 39] and references therein, using asymptotic reflection couplings. Additionally, in the symmetric case, i.e. the invariant probability measure has explicit representation, exponential ergodicity in the $L^{2}$-Wasserstein distance has been proven under the condition that the drift is bounded in the distribution variable and the disffusion is non-degenerate and only depends on the distribution variable, see [45]. For recent progresses on long-time behaviour of McKean-Vlasov SDEs, please refer to [43] on exponential ergodicity in total variation distance, [14] concerning exponential ergodicity in $L^{2}$-Wasserstein distance and mean field entropy for kinetic McKean-Vlasov SDEs, and [26] regarding non-asymptotic entropic bounds of stochastic algorithms.
Our main results show exponential ergodicity in both the $L^{2}$-Wasserstein distance and relative entropy for partially dissipative drifts, possibly unbounded in the distribution variable, with non-degenerate or degenerate diffusion. This further expands the results in [30] towards partially dissipative drifts and the results in [4, 32] towards drifts, which are not of gradient-type.
As a main tool in the non-degenerate case, we will use Wang’s observation in [41], which states that exponential ergodicity in relative entropy can be deduced from log-Sobolev inequalities. The study of log-Sobolev inequalities originated in the works of Leonard Gross [12], who introduced this for Gaussian measures. The seminal work by Bakry and Émery [1] links the log-Sobolev inequality to lower curvature bounds of diffusion processes via the so-called $\Gamma$-calculus introduced therein. Since then, log-Sobolev inequalities are an integral part in the study of SDEs and PDEs, e.g. in [6, 10, 13, 20, 22, 27, 33].
In what follows, we will first provide the log-Sobolev inequality for the time-marginal distribution of the decoupled SDE, given by
(1.2) $\text{\rm{d}}X_{s,t}^{\mu,x}=b_{t}(X_{s,t}^{\mu,x},\mu_{t})\text{\rm{d}}t+\sigma_{t}(X_{s,t}^{\mu,x},\mu_{t})\text{\rm{d}}W_{t},\ \ X_{s,s}^{\mu,x}=x\in\mathbb{R}^{d},t\geq s\geq 0,$
for any $\mu_{\cdot}\in C([0,\infty),\mathscr{P}_{2}(\mathbb{R}^{d}))$, i.e. we will prove the log-Sobolev inequality for the associated time-inhomogeneous semigroup defined by
(1.3) $P_{s,t}^{\mu}f(x)=\mathbb{E}f(X_{s,t}^{\mu,x}),\ \ t\geq s\geq 0,f\in\mathscr{B}_{b}(\mathbb{R}^{d}).$
These results apply to both the non-degenerate and degenerate case.
In order to obtain the log-Sobolev inequality for the invariant probability measure, we focus on the time-homogeneous case and use the hyperboundedness derived from Wang’s Harnack inequality. This improves the results in [30] towards partially dissipative drifts and the results in [32, 4] towards drifts, which are not of gradient-type.
The remainder of this paper is organized as follows. In Section 2, we study the non-degenerate case and we prove the log-Sobolev inequality for $P_{s,t}^{\mu}$ and $P_{t}^{\ast}$ (Theorem 2.1) under the monotonicity assumption (A). In the partially dissipative setting (H), we derive the log-Sobolev inequality for the invariant probability measure (Theorem 2.6). As an application, we deduce the exponential ergodicity in $\mathbb{W}_{2}$ and relative entropy in Theorem 2.7. In Section 3, we concentrate on the degenerate case. We establish the log-Sobolev inequality for $P_{s,t}^{\mu}$ and $P_{t}^{\ast}$ under the Lipschitz condition (B). For partially dissipative drifts satisfying ((C)), we derive the log-Sobolev inequality for the invariant probability measure in Theorem 3.5. As an application, we obtain exponential ergodicity in $\mathbb{W}_{2}$ and relative entropy for both uniformly dissipative (Theorem 3.7) and partially dissipative drifts (Theorem 3.9).
2 Non-degenerate case
2.1 Log-Sobolev inequality for $P_{s,t}^{\mu}$ and $P_{t}^{\ast}\mu_{0}$
We will adopt the $\Gamma$-calculus developed in [1] to our SDE setting. To this end, for any $\mu_{\cdot}\in C([0,\infty),\mathscr{P}_{2}(\mathbb{R}^{d}))$, we define
$b_{t}^{\mu}(x)=b_{t}(x,\mu_{t}),\ \ \sigma_{t}^{\mu}(x)=\sigma_{t}(x,\mu_{t}).$
The associated diffusion operator is then given by
(2.1) $\mathscr{L}_{t}^{\mu}f=\frac{1}{2}\mathrm{tr}(\sigma_{t}^{\mu}(\sigma_{t}^{\mu})^{\ast}\nabla^{2}f)+\nabla_{b_{t}^{\mu}}f,\ \ f\in C_{0}^{\infty}(\mathbb{R}^{d}),$
which gives rise to the $\Gamma$-operators
$\Gamma^{t,\mu}_{1}(f,g)=$ $\mathscr{L}_{t}^{\mu}(fg)-g\mathscr{L}_{t}^{\mu}f-f\mathscr{L}_{t}^{\mu}g=\langle(\sigma_{t}^{\mu})^{\ast}\nabla f,(\sigma_{t}^{\mu})^{\ast}\nabla g\rangle,$
$\Gamma^{t,\mu}_{2}(f,g)=$ $\mathscr{L}_{t}^{\mu}(\Gamma_{1}^{t,\mu}(f,g))-\Gamma_{1}^{t,\mu}(\mathscr{L}_{t}^{\mu}f,g)-\Gamma_{1}^{t,\mu}(\mathscr{L}_{t}^{\mu}g,f).$
For simplicity, we denote $\Gamma^{t,\mu}_{i}(f)=\Gamma^{t,\mu}_{i}(f,f),i=1,2$.
We make the following assumptions:
  1. (A)
    For any $t\geq 0$, $b_{t}$ and $\sigma_{t}$ are continuous in $\mathbb{R}^{d}\times\mathscr{P}_{2}(\mathbb{R}^{d})$. There exist constants $K_{1},K_{I}>0$ such that for all $t\geq 0$, $x,y\in\mathbb{R}^{d}$ and $\gamma,\tilde{\gamma}\in\mathscr{P}_{2}(\mathbb{R}^{d})$,
    $2\langle b_{t}(x,\gamma)-b_{t}(y,\tilde{\gamma}),x-y\rangle+\|\sigma_{t}(x,\gamma)-\sigma_{t}(y,\tilde{\gamma})\|_{HS}^{2}$
    (2.2) $\leq K_{1}|x-y|^{2}+K_{I}\mathbb{W}_{2}(\gamma,\tilde{\gamma})^{2}.$
Here, $K_{I}$ represents the dependence of $b$ and $\sigma$ on measure variable.
Under (A), (1.1) is well-posed in $\mathscr{P}_{2}(\mathbb{R}^{d})$ and recall that $P_{t}^{\ast}\gamma$ is the distribution of the solution to (1.1) with initial distribution $\gamma\in\mathscr{P}_{2}(\mathbb{R}^{d})$. Moreover, for any $\mu_{\cdot}\in C([0,\infty),\mathscr{P}_{2}(\mathbb{R}^{d}))$, (1.2) has a unique solution $X_{s,t}^{\mu,x}$ with associated time-inhomogenous semigroup $P_{s,t}^{\mu}$. See e.g. [39, Theorem 3.3.1] for these facts.
Theorem 2.1.
Assume $({\bf A})$ and suppose $\sigma_{t}(x,\gamma)=\sigma_{t}(\gamma)$ satisfying for some constant $\delta_{1}>0$,
(2.3) $\sigma_{t}(\gamma)\sigma_{t}(\gamma)^{\ast}\leq\delta_{1}I_{d\times d},\ \ \gamma\in\mathscr{P}_{2}(\mathbb{R}^{d}).$
Then, $P_{s,t}^{\mu}$ satisfies the log-Sobolev inequality: for any $\mu_{\cdot}\in C([0,\infty),\mathscr{P}_{2}(\mathbb{R}^{d}))$ and $0<f\in C_{0}^{\infty}(\mathbb{R}^{d})$, the smooth functions with compact support on $\mathbb{R}^{d}$,
(2.4) $P_{s,t}^{\mu}(f\log f)-P_{s,t}^{\mu}(f)\log P_{s,t}^{\mu}(f)\leq\frac{2\delta_{1}}{K_{1}}(\text{\rm{e}}^{K_{1}(t-s)}-1)P_{s,t}^{\mu}\big{(}|\nabla f^{\frac{1}{2}}|^{2}\big{)},\ \ 0\leq s\leq t.$
If moreover $\mu_{0}$ satisfies the log-Sobolev inequality: for some $L_{0}\geq 0$,
[…]
then $P_{t}^{\ast}\mu_{0}$ satisfies the log-Sobolev inequality:
(2.6) $\begin{split}(P_{t}^{\ast}\mu_{0})&(f\log f)-(P_{t}^{\ast}\mu_{0})(f)\log(P_{t}^{\ast}\mu_{0})(f)\\ &\leq 2\Big{(}L_{0}\text{\rm{e}}^{K_{1}t}+\frac{\delta_{1}}{K_{1}}(\text{\rm{e}}^{K_{1}t}-1)\Big{)}(P_{t}^{\ast}\mu_{0})\big{(}|\nabla f^{\frac{1}{2}}|^{2}\big{)},\ \ 0<f\in C_{0}^{\infty}(\mathbb{R}^{d}),t\geq 0.\end{split}$
Proof.
(1) By an approximation trick, we may and do assume that $b_{t}(\cdot,\gamma)$ is smooth. Recall that $X_{s,t}^{\mu,x}$ and $P_{s,t}^{\mu}$ is defined in (1.2) and (1.3). By ((A)) and $\sigma_{t}(x,\gamma)=\sigma_{t}(\gamma)$, it is standard to derive that
$|X_{s,t}^{\mu,x}-X_{s,t}^{\mu,y}|\leq\text{\rm{e}}^{\frac{1}{2}K_{1}(t-s)}|x-y|,\ \ x,y\in\mathbb{R}^{d},0\leq s\leq t.$
This implies that for any $x\in\mathbb{R}^{d},0\leq s\leq t,f\in C_{0}^{\infty}(\mathbb{R}^{d})$,
(2.7) $|\nabla P_{s,t}^{\mu}f|(x)$ $:=\limsup_{y\rightarrow x}\frac{|P_{s,t}^{\mu}f(y)-P_{s,t}^{\mu}f(x)|}{|y-x|}\leq\text{\rm{e}}^{\frac{1}{2}K_{1}(t-s)}P_{s,t}^{\mu}|\nabla f|(x).$
Since $\mathscr{L}_{t}^{\mu}$ given in (2.1) is a diffusion operator (see e.g. [3, Definition 1.11.1]), we have
$\mathscr{L}_{t}^{\mu}\{f\log f\}$ $=\frac{1}{2}f^{-1}\Gamma_{1}^{t,\mu}(f)+(\log f+1)\mathscr{L}_{t}^{\mu}f,\ \ f\in C_{0}^{\infty}(\mathbb{R}^{d}),f>0.$
Thus, Itô’s formula and the Kolmogorov backward equation
$\frac{\text{\rm{d}}P_{r,t}^{\mu}f}{\text{\rm{d}}r}=-\mathscr{L}_{r}^{\mu}P_{r,t}^{\mu}f,\ \ f\in C_{0}^{\infty}(\mathbb{R}^{d})$
imply that for $f\in C_{0}^{\infty}(\mathbb{R}^{d})$ with $f>0$,
$\text{\rm{d}}\{(P_{r,t}^{\mu}f)\log(P_{r,t}^{\mu}f)\}(X_{s,r}^{\mu,x})$ $=(-(\log(P_{r,t}^{\mu}f)+1)\mathscr{L}_{r}^{\mu}P_{r,t}^{\mu}f)(X_{s,r}^{\mu,x})\text{\rm{d}}r$
$\quad+\mathscr{L}_{r}^{\mu}\{(P_{r,t}^{\mu}f)\log(P_{r,t}^{\mu}f)\}(X_{s,r}^{\mu,x})\text{\rm{d}}r+\text{\rm{d}}M_{r}$
$=\frac{1}{2}[(P_{r,t}^{\mu}f)^{-1}[\Gamma_{1}^{r,\mu}(P_{r,t}^{\mu}f)]](X_{s,r}^{\mu,x})\text{\rm{d}}r+\text{\rm{d}}M_{r},\ \ r\in[s,t]$
for some martingale $M_{r}$. Integrating from $s$ to $t$ and taking expectation, we deduce from (2.3) and (2.7) that
$\begin{split}&P_{s,t}^{\mu}(f\log f)(x)-P_{s,t}^{\mu}(f)(x)\log P_{s,t}^{\mu}(f)(x)\\ &=\frac{1}{2}\int_{s}^{t}\mathbb{E}[(P_{r,t}^{\mu}f)^{-1}[\Gamma_{1}^{r,\mu}(P_{r,t}^{\mu}f)]](X_{s,r}^{\mu,x})\text{\rm{d}}r\\ &\leq\frac{1}{2}\delta_{1}\int_{s}^{t}\mathbb{E}[(P_{r,t}^{\mu}f)^{-1}|\nabla P_{r,t}^{\mu}f|^{2}](X_{s,r}^{\mu,x})\text{\rm{d}}r\\ &\leq\frac{1}{2}\delta_{1}\int_{s}^{t}\mathbb{E}[(P_{r,t}^{\mu}f)^{-1}\text{\rm{e}}^{K_{1}(t-r)}(P_{r,t}^{\mu}|\nabla f|)^{2}](X_{s,r}^{\mu,x})\text{\rm{d}}r.\end{split}$
This, combining with the following fact that
(2.8) $(P_{r,t}^{\mu}|\nabla f|)^{2}=4\big{(}P_{r,t}^{\mu}\big{(}|\nabla f^{\frac{1}{2}}|\sqrt{f}\big{)}\big{)}^{2}\leq 4P_{r,t}^{\mu}|\nabla f^{\frac{1}{2}}|^{2}P_{r,t}^{\mu}f,$
and the semigroup property $P_{s,r}^{\mu}P_{r,t}^{\mu}=P_{s,t}^{\mu}$, yields (2.4).
(2) Inserting $\mu_{t}=P_{t}^{\ast}\mu_{0},t\geq 0$ for $\mu_{0}\in\mathscr{P}_{2}(\mathbb{R}^{d})$, we obtain
(2.9) $(P_{t}^{\ast}\mu_{0})(f):=\int_{\mathbb{R}^{d}}f(x)(P_{t}^{\ast}\mu_{0})(\text{\rm{d}}x)=\int_{\mathbb{R}^{d}}P_{t}^{\mu}f(x)\mu_{0}(\text{\rm{d}}x),\ \ f\in\mathscr{B}_{b}(\mathbb{R}^{d}),$
where $P_{t}^{\mu}:={P}_{0,t}^{\mu}$. Integrating with respect to $\mu_{0}$ in (2.4) with $s=0$ and applying (2.9), we derive for $0<f\in C_{0}^{\infty}(\mathbb{R}^{d})$,
(2.10) $(P_{t}^{\ast}\mu_{0})(f\log f)-\mu_{0}(P_{t}^{\mu}(f)\log P_{t}^{\mu}(f))\leq 2\delta_{1}\frac{\text{\rm{e}}^{K_{1}t}-1}{K_{1}}(P_{t}^{\ast}\mu_{0})\left(|\nabla f^{\frac{1}{2}}|^{2}\right).$
By (2.5), (2.7) and (2.8) for $r=0$ and (2.9), we conclude for $0<f\in C_{0}^{\infty}(\mathbb{R}^{d})$,
$\mu_{0}(P_{t}^{\mu}(f)\log P_{t}^{\mu}(f))$ $\leq(P_{t}^{\ast}\mu_{0})(f)\log(P_{t}^{\ast}\mu_{0})(f)+2L_{0}\mu_{0}\big{(}|\nabla(\bar{P}_{t}^{\mu_{0}}f)^{\frac{1}{2}}|^{2}\big{)}$
$\leq(P_{t}^{\ast}\mu_{0})(f)\log(P_{t}^{\ast}\mu_{0})(f)+2L_{0}\text{\rm{e}}^{K_{1}t}(P_{t}^{\ast}\mu_{0})\big{(}|\nabla f^{\frac{1}{2}}|^{2}\big{)}.$
This, together with (2.10), implies (2.6) and the proof is completed. ∎
Remark 2.2.
When $\sigma=I_{d\times d}$, the assertions in Theorem 2.1 have been verified in [25, Proposition 1.1]. One can also refer to [8, Theorem 4.1] for the log-Sobolev inequality of time-inhomogeneous semigroup and [29, Lemma 5.2] for the reflecting case under the Bakry-Émery curvature condition. Moreover, (2.7) is equivalent to [8, (10)] if we assume $\sigma\sigma^{\ast}\geq\delta_{2}$ for some $\delta_{2}>0$.
Remark 2.3.
In [7], the authors derive functional inequalities for the heat flow on Riemannian manifolds with time-dependent metrics. More precisely, they consider the operator $L_{t}=\Delta_{t}+Z_{t}$ for some vector field $Z_{t}$ on a complete Remannian manifold $M$ with time–dependent Riemannian metric $g_{t}$. They use the following curvature condition:
$\mathfrak{R}_{t}^{Z}(X,X)$ $:=\mathrm{Ric}_{t}(X,X)-\langle\nabla_{X}^{t}Z_{t},X\rangle-\frac{1}{2}\partial_{t}g_{t}(X,X)\geq k(t),$
(2.11) $\ \ X\in T_{x}M\text{ such that }g_{t}(X,X)=1,x\in M,k\in C([0,\infty)\times M).$
When $\sigma$ depends on the spatial variable and $\sigma_{t}^{\mu}(\sigma_{t}^{\mu})^{\ast}$ is invertible, we may define a Riemannian metric as
$g_{t}(X,Y)=\langle(\sigma_{t}^{\mu}(\sigma_{t}^{\mu})^{\ast})^{-1}X,Y\rangle.$
Then the functional inequalities for $P_{s,t}^{\mu}$ follow by using the conclusions in [7] provided that (2.3) holds. However, as explained in [40, p.1451], the assumption (2.3) involves the term $\nabla^{2}[(\sigma_{t}^{\mu}(\sigma_{t}^{\mu})^{\ast})^{-1}]$, which is more or less implicit. This is the reason why we only consider the case $\sigma_{t}(x,\gamma)=\sigma_{t}(\gamma)$.
Remark 2.4.
If $\sigma=I_{d\times d}$ and $b$ is of gradient-type, then the $L^{1}$-gradient estimate (2.7), the log-Sobolev inequality for $P_{s,t}^{\mu}$ and the Bakry–Émery curvature condition
$\Gamma^{t,\mu}_{2}(f)\geq\rho\Gamma_{1}^{t,\mu}(f)=\rho|\nabla f|^{2}$
for some constant $\rho\in\mathbb{R}$ are equivalent, see e.g. [22, Proposition 3.12].
2.2 Log-Sobolev inequality for the invariant probability measure of time-homogeneous decoupled SDEs
From now on we will focus on the time-homogeneous version of (1.1), i.e.,
(2.12) $\text{\rm{d}}X_{t}=b(X_{t},\mathscr{L}_{X_{t}})\text{\rm{d}}t+\sigma(X_{t},\mathscr{L}_{X_{t}})\text{\rm{d}}W_{t},$
and the associated time-homogeneous decoupled SDE, i.e. for any $\mu\in\mathscr{P}_{2}(\mathbb{R}^{d})$, consider
(2.13) $\text{\rm{d}}\tilde{X}_{t}^{\mu}=b(\tilde{X}_{t}^{\mu},\mu)\text{\rm{d}}t+\sigma(\tilde{X}_{t}^{\mu},\mu)\text{\rm{d}}W_{t}.$
Let $(\tilde{P}_{t}^{\mu})^{\ast}\nu$ be the distribution of the solution to (2.13) with initial distribution $\nu\in\mathscr{P}_{2}(\mathbb{R}^{d})$ and associated time-homogeneous semigroup
$\tilde{P}_{t}^{\mu}f(x)=\int_{\mathbb{R}^{d}}f\text{\rm{d}}\{(\tilde{P}_{t}^{\mu})^{\ast}\delta_{x}\},\ \ f\in\mathscr{B}_{b}(\mathbb{R}^{d}),x\in\mathbb{R}^{d},t\geq 0.$
To establish the log-Sobolev inequality for the invariant probability measure of (2.13), we make use of the following partially dissipative condition.
  1. (H)
    $b$ and $\sigma$ are continuous in $\mathbb{R}^{d}\times\mathscr{P}_{2}(\mathbb{R}^{d})$. There exist constants $K_{1}>0,K_{2}>0,K_{I}>0,r_{0}>0$ and $\delta_{1}\geq\delta_{2}>0$ such that for any $x,y\in\mathbb{R}^{d},\gamma,\tilde{\gamma}\in\mathscr{P}_{2}(\mathbb{R}^{d})$,
    $2\langle b(x,\gamma)-b(y,\tilde{\gamma}),x-y\rangle+\|\sigma(x,\gamma)-\sigma(y,\tilde{\gamma})\|_{HS}^{2}$
    (2.14) $\leq K_{1}|x-y|^{2}1_{\{|x-y|\leq r_{0}\}}-K_{2}|x-y|^{2}1_{\{|x-y|>r_{0}\}}+K_{I}\mathbb{W}_{2}(\gamma,\tilde{\gamma})^{2},$
    and
    (2.15) $\delta_{2}I_{d\times d}\leq\sigma\sigma^{\ast}\leq\delta_{1}I_{d\times d}.$
Before we prove our next main result Theorem 2.6, we make a few comments on the proof. First, by Wang’s Harnack inequality (2.2) for $\tilde{P}_{t}^{\mu}$, the hyperboundedness (2.26) for $\tilde{P}_{t}^{\mu}$ holds. From this we can employ Theorem 2.1(1) in [28], which provides the defective log-Sobolev inequality for the invariant probability measure (IPM) of (2.13). Let us remark, that although [28, Theorem 2.1(1)] requires a curvature lower bound, the proof of it reveals that only the log-Sobolev inequality of the semigroup $\tilde{P}_{t}^{\mu}$ is used.
In order to prove Theorem 2.6, we need to derive a defective log-Sobolev inequality for the IPM of (2.13) uniformly in $\hat{\mathscr{P}}_{2,M}(\mathbb{R}^{d}):=\{\mu\in\mathscr{P}_{2}(\mathbb{R}^{d}),\mu(|\cdot|^{2})\leq M\}$ for any $M>0$. Together with tightness and uniform bounds on the densities of each IPM, this implies the result, as the next lemma shows.
Lemma 2.5.
  1. (1)
    Let $L=\frac{1}{2}\mathrm{tr}(a\nabla^{2})+\langle b,\nabla\rangle$ be the generator of a conservative diffusion process on $\mathbb{R}^{d}$ with associated semigroup $(P_{t})_{t\geq 0}$. Let $\bar{\mu}\in\mathscr{P}(\mathbb{R}^{d})$ be the IPM of $(P_{t})_{t\geq 0}$. Assume that there exist $t_{0}>0$, $1<p<q<\infty$ and $C_{0}>0$ such that $\|P_{t_{0}}\|_{p\rightarrow q}:=\|P_{t_{0}}\|_{L^{p}(\bar{\mu})\rightarrow L^{q}(\bar{\mu})}<C_{0}.$ If additionally the log-Sobolev inequality for $P_{t}$ holds, i.e.
    $P_{t}(f\log f)-P_{t}f\log P_{t}f\leq 4h(t)P_{t}(|\nabla f^{\frac{1}{2}}|^{2}),\ \ t\geq 0,0<f\in C_{0}^{\infty}(\mathbb{R}^{d}),$
    for some measurable function $h:[0,\infty)\rightarrow[0,\infty)$, then $\bar{\mu}$ satisfies the following defective log-Sobolev inequality:
    $\bar{\mu}(f\log f)-\bar{\mu}(f)\log\bar{\mu}(f)$
    $\leq\frac{4p(q-1)}{q-p}h(t_{0})\bar{\mu}(|\nabla f^{\frac{1}{2}}|^{2})+\frac{pq}{q-p}\log\|P_{t_{0}}\|_{p\rightarrow q},\ \ \bar{\mu}(f)=1,0<f\in C_{0}^{\infty}(\mathbb{R}^{d}).$
  2. (2)
    Let $L^{\theta}=\frac{1}{2}\mathrm{tr}(a^{\theta}\nabla^{2})+\langle b^{\theta},\nabla\rangle$, $\theta\in\Theta$ be a family of conservative diffusion processes on $\mathbb{R}^{d}$ with associated semigroup $(P_{t}^{\theta})_{t\geq 0}$ satisfying the assumptions in (1) uniformly in $\theta$. Assume that the IPM $\bar{\mu}^{\theta}$ of $(P_{t}^{\theta})_{t\geq 0}$ form a tight subset of $\mathscr{P}(\mathbb{R}^{d})$ and that for any $R>0$, there exist positive constants $C_{1}(R)\leq C_{2}(R)$ such that $\bar{\rho}_{\theta}=\frac{\text{\rm{d}}\bar{\mu}^{\theta}}{\text{\rm{d}}x},\ \ \theta\in\Theta$ satisfy
    (2.16) $C_{1}(R)\leq\bar{\rho}_{\theta}(x)\leq C_{2}(R),\ \ x\in B_{R}:=\{x\in\mathbb{R}^{d},|x|\leq R\},\theta\in\Theta.$
    Then the weak Poincaré inequality holds, i.e.
    $\bar{\mu}^{\theta}(f^{2})\leq\alpha(r)\bar{\mu}^{\theta}(|\nabla f|^{2})+r\|f\|_{\infty}^{2},\ \ r>0,0<f\in C_{0}^{\infty}(\mathbb{R}^{d}),\bar{\mu}^{\theta}(f)=0,\theta\in\Theta$
    for some decreasing function $\alpha:(0,\infty)\rightarrow[0,\infty).$ Moreover, the log-Sobolev inequality holds, that is there exists a constant $C_{LS}>0$ such that
[…]
Proof.
(1) follows from he proof of [28, Theorem 2.1(1)] so that we omit it. It remains to prove (2). Observe that for any $R>0$, the classical Poincaré inequality for Lebesgue measure holds, i.e.
(2.18) $\frac{\int_{B_{R}}f^{2}\text{\rm{d}}x}{|B_{R}|}\leq\frac{\left(\int_{B_{R}}f\text{\rm{d}}x\right)^{2}}{|B_{R}|^{2}}+\frac{\beta(R)}{|B_{R}|}\int_{B_{R}}|\nabla f|^{2}\text{\rm{d}}x,\ \ f\in C_{0}^{\infty}(\mathbb{R}^{d})$
for some constant $\beta(R)>0$. When $(\bar{\mu}^{\theta})_{\theta\in\Theta}$ is tight, for any $\varepsilon\in(0,1)$, we can take $R>0$ large enough such that $\bar{\mu}^{\theta}(B_{R})\geq 1-\varepsilon,\theta\in\Theta$. Let $\text{\rm{d}}\bar{\mu}^{\theta}_{R}=\frac{1_{B_{R}}}{\bar{\mu}^{\theta}(B_{R})}\text{\rm{d}}\bar{\mu}^{\theta}$. It follows from (2.16) and (2.18) that
$\int_{\mathbb{R}^{d}}(f-\bar{\mu}^{\theta}_{R}(f))^{2}\text{\rm{d}}\bar{\mu}^{\theta}_{R}$
$=\inf_{c\in\mathbb{R}}\int_{\mathbb{R}^{d}}(f-c)^{2}\text{\rm{d}}\bar{\mu}^{\theta}_{R}$
$=\inf_{c\in\mathbb{R}}\frac{|B_{R}|}{\bar{\mu}^{\theta}(B_{R})}\int_{\mathbb{R}^{d}}(f-c)^{2}\frac{1_{B_{R}}}{|B_{R}|}\rho\text{\rm{d}}x$
$\leq\frac{\beta(R)C_{2}(R)}{\bar{\mu}^{\theta}(B_{R})}\int_{\mathbb{R}^{d}}1_{B_{R}}|\nabla f|^{2}\text{\rm{d}}x$
$\leq\frac{\beta(R)C_{2}(R)}{C_{1}(R)\bar{\mu}^{\theta}(B_{R})}\int_{\mathbb{R}^{d}}|\nabla f|^{2}\text{\rm{d}}\bar{\mu}^{\theta},\ \ f\in C_{0}^{\infty}(\mathbb{R}^{d}),\theta\in\Theta.$
This means
$\int_{B_{R}}f^{2}\text{\rm{d}}\bar{\mu}^{\theta}\leq\frac{\left(\int_{B_{R}}f\text{\rm{d}}\bar{\mu}^{\theta}\right)^{2}}{\bar{\mu}^{\theta}(B_{R})}+\frac{\beta(R)C_{2}(R)}{C_{1}(R)}\int_{\mathbb{R}^{d}}|\nabla f|^{2}\text{\rm{d}}\bar{\mu}^{\theta},\ \ f\in C_{0}^{\infty}(\mathbb{R}^{d}),\theta\in\Theta.$
This implies that
$\bar{\mu}^{\theta}(f^{2})\leq\frac{\beta(R)C_{2}(R)}{C_{1}(R)}\bar{\mu}^{\theta}(|\nabla f|^{2})+\frac{\varepsilon}{1-\varepsilon}\|f\|_{\infty}^{2},\ \ f\in C_{0}^{\infty}(\mathbb{R}^{d}),\bar{\mu}^{\theta}(f)=0,\theta\in\Theta.$
This together with the fact that $(\bar{\mu}^{\theta})_{\theta\in\Theta}$ is tight implies the weak Poincaré inequality
$\bar{\mu}^{\theta}(f^{2})\leq\alpha(r)\bar{\mu}^{\theta}(|\nabla f|^{2})+r\|f\|_{\infty}^{2},\ \ r>0,f\in C_{0}^{\infty}(\mathbb{R}^{d}),\bar{\mu}^{\theta}(f)=0,\theta\in\Theta$
holds for
$\alpha(r):=\inf\left\{\frac{\beta(R)C_{2}(R)}{C_{1}(R)}:\inf_{\theta\in\Theta}\bar{\mu}^{\theta}(B_{R})\geq\frac{1}{1+r}\right\}.$
We can take $r_{0}>0$ satisfying $c_{0}:=\frac{1}{2}(1+r_{0}+\sqrt{(C_{2}+1+r_{0})r_{0}})<1$ such that
(2.19) $\bar{\mu}^{\theta}(f^{2})\leq\alpha(r_{0})\bar{\mu}^{\theta}(|\nabla f|^{2})+r_{0}\|f\|_{\infty}^{2},\ \ f\in C_{0}^{\infty}(\mathbb{R}^{d}),\bar{\mu}^{\theta}(f)=0,\theta\in\Theta.$
Observe that due to [38, Theorem 3.3.1], (1) implies the defective Poincaré inequality
$\bar{\mu}^{\theta}(f^{2})\leq C_{1}\bar{\mu}^{\theta}(|\nabla f|^{2})+C_{2}\bar{\mu}^{\theta}(|f|)^{2},\ \ \theta\in\Theta.$
By [35, Proposition 4.1.2(2)], we conclude from this and (2.19) that the Poincaré inequality holds:
$\bar{\mu}^{\theta}(f^{2})-\bar{\mu}^{\theta}(f)^{2}\leq\frac{C_{1}+\alpha(r_{0})}{1-c_{0}}\bar{\mu}^{\theta}(|\nabla f|^{2}),\ \ f\in C_{0}^{\infty}(\mathbb{R}^{d}),\theta\in\Theta.$
This together with the defective log-Sobolev inequality implies (2.17), see for instance the proof of [37, Corollary 1.3]. ∎
Theorem 2.6.
Assume (H). Then the following assertions hold.
  1. $(1)$
    For any $\mu\in\mathscr{P}_{2}(\mathbb{R}^{d})$, (2.13) has a unique invariant probability measure denoted by $\Phi(\mu)$. Moreover, if $K_{I}<K_{2}$, there exists an $M>0$ such that $\Phi$ is a map on $\hat{\mathscr{P}}_{2,M}(\mathbb{R}^{d})$, where
    $\hat{\mathscr{P}}_{2,M}(\mathbb{R}^{d}):=\left\{\mu\in\mathscr{P}_{2}(\mathbb{R}^{d}):\mu(|\cdot|^{2})\leq M\right\}.$
    Moreover, there exists a constant $\varepsilon>0$ and $C_{1}>0$ such that
    (2.20) $\{\Phi(\mu)\}(\text{\rm{e}}^{\varepsilon|\cdot|^{2}})\leq C_{1},\ \ \mu\in\hat{\mathscr{P}}_{2,M}(\mathbb{R}^{d}).$
    Consequently, there exist constants $t_{0}>0,p>1,C_{2}>0$ such that
    $\|\tilde{P}_{t_{0}}^{\mu}\|_{L^{p}(\Phi(\mu))\rightarrow L^{2p}(\Phi(\mu))}\leq C_{2},\ \ \mu\in\hat{\mathscr{P}}_{2,M}(\mathbb{R}^{d}).$
  2. $(2)$
    If in addition, $\sigma(x,\gamma)=\sigma(\gamma)$, then $\Phi(\mu)$ satisfies the log-Sobolev inequality
    (2.21) $\Phi(\mu)(f^{2}\log f^{2})\leq C_{L}\Phi(\mu)(|\nabla f|^{2}),$
    for some constant $C_{L}>0,$ any $\mu\in\hat{\mathscr{P}}_{2,M}(\mathbb{R}^{d}),$ and $f\in C_{0}^{\infty}(\mathbb{R}^{d})$ with $\Phi(\mu)(f^{2})=1$.
Proof.
(1)(i) Taking $y={\bf 0}$ in ((H)), we get
$2\langle b(x,\mu)-b({\bf 0},\delta_{\bf 0}),x\rangle+\|\sigma(x,\mu)-\sigma({\bf 0},\delta_{\bf 0})\|_{HS}^{2}$
(2.22) $\leq K_{1}|x|^{2}1_{\{|x|\leq r_{0}\}}-K_{2}|x|^{2}1_{\{|x|>r_{0}\}}+K_{I}\|\mu\|_{2}^{2},\ \ x\in\mathbb{R}^{d},\mu\in\mathscr{P}_{2}(\mathbb{R}^{d})$
So, it follows from (2.15) and (2.2) that
(2.23) $\begin{split}&2\langle x,b(x,\mu)\rangle+\|\sigma(x,\mu)\|_{\rm HS}^{2}\\ &\leq-K_{2}|x|^{2}+(K_{1}+K_{2})r_{0}^{2}+2|x||b({\bf 0},\delta_{\bf 0})|+K_{I}\mu(|\cdot|^{2})\\ &\quad+\|\sigma({\bf 0},\delta_{\bf 0})\|_{\rm HS}^{2}+2\|\sigma(x,\mu)-\sigma({\bf 0},\delta_{\bf 0})\|_{HS}\|\sigma({\bf 0},\delta_{\bf 0})\|_{\rm HS}\\ &\leq-K_{2}|x|^{2}+2|x||b({\bf 0},\delta_{\bf 0})|+K_{I}\mu(|\cdot|^{2})+C_{0},\ \ x\in\mathbb{R}^{d},\mu\in\mathscr{P}_{2}(\mathbb{R}^{d})\end{split}$
for some constant $C_{0}>0$. Whence, an application of [44, Lemma 2.7], along with (2.15) and (2.23), yields that (2.13) has a unique IPM, written as $\Phi(\mu).$ Next, once $K_{2}>K_{I}$ is valid, [44, Lemma 2.8] implies that, for some constant $M>0,$ $\hat{\mathscr{P}}_{2,M}(\mathbb{R}^{d})$ is an invariant set of $\Phi$ so the second assertion in (1) holds true.
(ii) Itô’s formula, (2.15) as well as (2.23) imply that for $\varepsilon=\frac{K_{2}}{4\delta_{1}}$ and $\mu\in\hat{\mathscr{P}}_{2,M}(\mathbb{R}^{d}),$
$\text{\rm{d}}\text{\rm{e}}^{\varepsilon|\tilde{X}_{t}^{\mu}|^{2}}$ $\leq\varepsilon\text{\rm{e}}^{\varepsilon|\tilde{X}_{t}^{\mu}|^{2}}\left\{-K_{2}|\tilde{X}_{t}^{\mu}|^{2}+2|\tilde{X}_{t}^{\mu}||b({\bf 0},\delta_{\bf 0})|+K_{I}\mu(|\cdot|^{2})+C_{0}+2\varepsilon\delta_{1}|\tilde{X}_{t}^{\mu}|^{2}\right\}\text{\rm{d}}t$
(2.24) $\quad+2\varepsilon\text{\rm{e}}^{\varepsilon|\tilde{X}_{t}^{\mu}|^{2}}\langle\sigma(\tilde{X}_{t}^{\mu},\mu)\text{\rm{d}}W_{t},\tilde{X}_{t}^{\mu}\rangle$
$\leq$ $\varepsilon\text{\rm{e}}^{\varepsilon|\tilde{X}_{t}^{\mu}|^{2}}\left\{-\frac{K_{2}}{4}|\tilde{X}_{t}^{\mu}|^{2}+\frac{4|b({\bf 0},\delta_{\bf 0})|^{2}}{K_{2}}+K_{I}M+C_{0}\right\}\text{\rm{d}}t$
$+2\varepsilon\text{\rm{e}}^{\varepsilon|\tilde{X}_{t}^{\mu}|^{2}}\langle\sigma(\tilde{X}_{t}^{\mu},\mu)\text{\rm{d}}W_{t},\tilde{X}_{t}^{\mu}\rangle.$
Note that we can find constants $c_{1}(\varepsilon),c_{2}(\varepsilon)>0$ such that
$\varepsilon\text{\rm{e}}^{\varepsilon|\tilde{X}_{t}^{\mu}|^{2}}\left\{-\frac{K_{2}}{4}|\tilde{X}_{t}^{\mu}|^{2}+\frac{4|b({\bf 0},\delta_{\bf 0})|^{2}}{K_{2}}+K_{I}M+C_{0}\right\}\leq c_{1}(\varepsilon)-c_{2}(\varepsilon)\text{\rm{e}}^{\varepsilon|\tilde{X}_{t}^{\mu}|^{2}}.$
Substituting this into (2.2) and taking $\tilde{X}_{0}^{\mu}={\bf 0}$ enables us to derive that for some constant $C_{1}>0$
$\int_{0}^{t}\mathbb{E}(\text{\rm{e}}^{\varepsilon|\tilde{X}_{s}^{\mu}|^{2}})\,\text{\rm{d}}s\leq C_{1}t.$
This obviously implies (2.20) by a tightness argument.
(iii) By [42, Lemma 2.6], there exist constants $p>1$ depending on $\delta_{1},\delta_{2}$, $h_{1}>0$ depending on $K_{1},r_{0}$, $\delta_{1},\delta_{2}$ and $h_{2}>0$ depending on $\delta_{1},\delta_{2}$ such that the dimension free Harnack inequality holds, i.e. for any $t>0$,
$(\tilde{P}_{t}^{\mu}|f|)^{p}(x)$ $\leq(\tilde{P}_{t}^{\mu}|f|^{p})(y)$
(2.25) ×$\qquad\quad\times\exp\left\{h_{1}t+\frac{h_{2}K_{2}|x-y|^{2}}{\text{\rm{e}}^{K_{2}t}-1}\right\},\ \ \mu\in\mathscr{P}_{2}(\mathbb{R}^{d}),x,y\in\mathbb{R}^{d},f\in\mathscr{B}_{b}(\mathbb{R}^{d}).$
When $\sigma=I_{d\times d}$, the Harnack inequality above has been established in [41, p.1138]. By following the strategy adopted in [40, Corollary 1.3], it is more or less standard to derive from (2.2) that for some locally bounded function $(\infty,0)\ni t\mapsto C_{t}\in[0,\infty)$,
$\sup_{\{\Phi(\mu)\}(|f|^{p})\leq 1}(\tilde{P}_{t}^{\mu}|f|)^{2p}(x)$ $\leq C_{t}\exp\left\{\frac{h_{2}K_{2}}{\text{\rm{e}}^{K_{2}t}-1}|x|^{2}\right\},\ \ \mu\in\mathscr{P}_{2}(\mathbb{R}^{d}),x\in\mathbb{R}^{d}.$
Then, by taking $t_{0}>0$ such that $\frac{h_{2}K_{2}}{\text{\rm{e}}^{K_{2}t_{0}}-1}=\varepsilon$, (2.20) enables us to derive that
(2.26) $\|\tilde{P}_{t_{0}}^{\mu}\|_{L^{p}(\Phi(\mu))\rightarrow L^{2p}(\Phi(\mu))}^{2p}\leq C_{t_{0}}\{\Phi(\mu)\}(\text{\rm{e}}^{\varepsilon|\cdot|^{2}})\leq C_{t_{0}}C_{1},\ \ \mu\in\hat{\mathscr{P}}_{2,M}(\mathbb{R}^{d}).$
(2) If $\sigma(x,\gamma)=\sigma(\gamma)$, by (2.4) with $s=0$, we have
$\tilde{P}_{t}^{\mu}(f\log f)-\tilde{P}_{t}^{\mu}(f)\log\tilde{P}_{t}^{\mu}(f)$
(2.27) $\leq\frac{2\delta_{1}}{K_{1}}(\text{\rm{e}}^{K_{1}t}-1)\tilde{P}_{t}^{\mu}\big{(}|\nabla f^{\frac{1}{2}}|^{2}\big{)},\ \ t\geq 0,\mu\in\hat{\mathscr{P}}_{2,M}(\mathbb{R}^{d}),0<f\in C_{0}^{\infty}(\mathbb{R}^{d}).$
This, together with (2.26) and Lemma 2.5(1), implies that the following defective log-Sobolev inequality
$\Phi(\mu)(f\log f)-\Phi(\mu)(f)\log\Phi(\mu)(f)$
$\leq\frac{2\delta_{1}}{K_{1}}(\text{\rm{e}}^{K_{1}t_{0}}-1)(2p-1)\Phi(\mu)(|\nabla f^{\frac{1}{2}}|^{2})+\log(C_{t_{0}}C_{1}),\ \ \mu\in\hat{\mathscr{P}}_{2,M}(\mathbb{R}^{d}),0<f\in C_{0}^{\infty}(\mathbb{R}^{d}).$
Since $\{\Phi(\mu):\mu\in\hat{\mathscr{P}}_{2,M}(\mathbb{R}^{d})\}$ is tight due to (2.20), according to Lemma 2.5(2), in order to derive the desired log-Sobolev inequality (2.21), it remains to prove that for any $R>0$ and $M>0$, there exist positive constants $C_{1}$ and $C_{2}$ depending on $R$, $M$ such that
(2.28) $C_{1}\leq\rho_{\mu}(x)\leq C_{2},\ \ x\in B_{R},\mu\in\hat{\mathscr{P}}_{2,M}(\mathbb{R}^{d})$
for $\rho_{\mu}:=\frac{\text{\rm{d}}\Phi(\mu)}{\text{\rm{d}}x}$. Since $b(x,\mu)$ is locally bounded in $(x,\mu)\in\mathbb{R}^{d}\times\mathscr{P}_{2}(\mathbb{R}^{d})$, for any $R>0,M>0$, there exists $\eta(R,M)>0$ such that
$\sup_{|x|\leq R}|b(x,\mu)|\leq\eta(R,M),\ \ \mu\in\hat{\mathscr{P}}_{2,M}(\mathbb{R}^{d}).$
Noting $(\sigma\sigma^{\ast})(\mu)\geq\delta_{2}I_{d\times d}$, by [5, Corollary 1.2.11], for any $R>0$ and $M>0$, there exists a constant $C_{0}$ depending on $R$, $\eta(R,M)$ and $\delta_{2}$ such that
$\sup_{\bar{B}_{R}}\rho_{\mu}\leq C_{0}\inf_{\bar{B}_{R}}\rho_{\mu},\ \ \mu\in\hat{\mathscr{P}}_{2,M}(\mathbb{R}^{d}).$
Take $R>0$ such that $\inf_{\mu\in\hat{\mathscr{P}}_{2,M}(\mathbb{R}^{d})}\Phi(\mu)(\bar{B}_{R})\geq\frac{1}{2}$, which can be ensured by (2.20). Then for any $x,y\in\bar{B}_{R}$, it holds
$\rho_{\mu}(x)\leq C_{0}\rho_{\mu}(y),\ \ \mu\in\hat{\mathscr{P}}_{2,M}(\mathbb{R}^{d}).$
This implies
$\frac{1}{2}\leq\Phi(\mu)(\bar{B}_{R})=\int_{\bar{B}_{R}}\rho_{\mu}(x)\text{\rm{d}}x\leq|\bar{B}_{R}|C_{0}\rho_{\mu}(y),\ \ y\in\bar{B}_{R},\mu\in\hat{\mathscr{P}}_{2,M}(\mathbb{R}^{d}),$
and
$\rho_{\mu}(x)|\bar{B}_{R}|\leq\int_{\bar{B}_{R}}C_{0}\rho_{\mu}(y)\text{\rm{d}}y\leq C_{0},\ \ x\in\bar{B}_{R},\mu\in\hat{\mathscr{P}}_{2,M}(\mathbb{R}^{d}).$
Therefore, (2.28) holds. ∎
2.3 Exponential ergodicity of (2.12) in $\mathbb{W}_{2}$ and relative entropy under partially dissipative condition
Theorem 2.7.
Assume (H) and that $\sigma(x,\gamma)=\sigma(\gamma)$. If $K_{I}$ is small enough, then (2.12) has a unique IPM $\mu_{\infty}$ which satisfies the log-Sobolev inequality
(2.29) $\mu_{\infty}(f^{2}\log f^{2})\leq C_{L}\mu_{\infty}(|\nabla f|^{2}),\ \ \mu_{\infty}(f^{2})=1,\ \ f\in C_{0}^{\infty}(\mathbb{R}^{d})$
for some constant $C_{L}>0$. Moreover, there exist constants $c>0,\lambda>0$ such that
(2.30) $\mathbb{W}_{2}(P_{t}^{\ast}\mu_{0},\mu_{\infty})\leq c\text{\rm{e}}^{-\lambda t}\mathbb{W}_{2}(\mu_{0},\mu_{\infty}),\ \ t\geq 0,\mu_{0}\in\mathscr{P}_{2}(\mathbb{R}^{d}).$
If in addition, $\sigma(x,\gamma)=\sigma$ and
$\langle b(x,\gamma)-b(y,\tilde{\gamma}),x-y\rangle$
(2.31) $\leq K|x-y|^{2}+K|x-y|\mathbb{W}_{2}(\gamma,\tilde{\gamma}),\ \ x,y\in\mathbb{R}^{d},\gamma,\tilde{\gamma}\in\mathscr{P}_{2}(\mathbb{R}^{d})$
for some constant $K\geq 0$, then it holds
(2.32) $\mathrm{Ent}(P_{t}^{\ast}\mu_{0}|\mu_{\infty})\leq\bar{c}^{2}\text{\rm{e}}^{-2\lambda t}\mathrm{Ent}(\mu_{0}|\mu_{\infty}),\ \ t\geq 1,\mu_{0}\in\mathscr{P}_{2}(\mathbb{R}^{d})$
for some constant $\bar{c}>0$.
Proof.
(1) Recall $\Phi(\mu)$ is introduced in Theorem 2.6. Combining (2.21) and [41, Theorem 2.6(2)], there exist constants $c_{1},\lambda_{1}>0$ independent of $\mu$ such that
(2.33) $\mathbb{W}_{2}((\tilde{P}_{t}^{\mu})^{\ast}\nu,\Phi(\mu))\leq c_{1}\text{\rm{e}}^{-\lambda_{1}t}\mathbb{W}_{2}(\nu,\Phi(\mu)),\ \ \mu\in\hat{\mathscr{P}}_{2,M}(\mathbb{R}^{d}),\nu\in\mathscr{P}_{2}(\mathbb{R}^{d}),t\geq 0.$
Define
$\delta_{0}:=\left(\inf_{t>\frac{\log c_{1}}{\lambda_{1}}}\frac{\sqrt{\frac{1}{K_{1}}(\text{\rm{e}}^{K_{1}t}-1)}}{1-c_{1}\text{\rm{e}}^{-\lambda_{1}t}}\right)^{-2},$
By taking advantage of [45, Proposition 2.2] and ((H)), when $K_{I}<\min(K_{2},\delta_{0})$, (2.12) has a unique invariant probability measure $\mu_{\infty}\in\hat{\mathscr{P}}_{2,M}(\mathbb{R}^{d})$. For reader’s convenience, we give more details. For $\mu_{1},\mu_{2}\in\mathscr{P}_{2}(\mathbb{R}^{d})$, let $\tilde{X}_{t}^{i},i=1,2$ solve
$\text{\rm{d}}\tilde{X}_{t}^{i}=b(\tilde{X}_{t}^{i},\mu_{i})\text{\rm{d}}t+\sigma(\mu_{i})\text{\rm{d}}W_{t}$
with $\tilde{X}_{0}^{1}=\tilde{X}_{0}^{2}$ satisfying $\mathscr{L}_{\tilde{X}_{0}^{1}}=\Phi(\mu_{1})$. Then it holds $\mathscr{L}_{\tilde{X}_{t}^{1}}=\Phi(\mu_{1})$ and $\mathscr{L}_{\tilde{X}_{t}^{2}}=(\tilde{P}_{t}^{\mu_{2}})^{\ast}(\Phi(\mu_{1}))$. It follows from ((H)) and Gronwall’s inequality that
(2.34) $\mathbb{W}_{2}((\tilde{P}_{t}^{\mu_{2}})^{\ast}(\Phi(\mu_{1})),\Phi(\mu_{1}))^{2}\leq\mathbb{E}|\tilde{X}_{t}^{1}-\tilde{X}_{t}^{2}|^{2}\leq\int_{0}^{t}K_{I}\text{\rm{e}}^{K_{1}(t-s)}\mathbb{W}_{2}(\mu_{1},\mu_{2})^{2}\text{\rm{d}}s,$
which together with (2.33) for $\mu=\mu_{2},\nu=\Phi(\mu_{1})$ implies
$\mathbb{W}_{2}(\Phi(\mu_{1}),\Phi(\mu_{2}))\leq c_{1}\text{\rm{e}}^{-\lambda_{1}t}\mathbb{W}_{2}(\Phi(\mu_{1}),\Phi(\mu_{2}))+\sqrt{K_{I}\frac{1}{K_{1}}(\text{\rm{e}}^{K_{1}t}-1)}\mathbb{W}_{2}(\mu_{1},\mu_{2}).$
Hence, combining this with Theorem 2.6 implies that when $K_{I}<\min\{K_{2},\delta_{0}\}$, $\Phi$ is a contractive map on $\hat{\mathscr{P}}_{2,M}(\mathbb{R}^{d})$ and has a unique fixed point $\mu_{\infty}$ satisfying $\Phi(\mu_{\infty})=\mu_{\infty}$, which means that (2.12) has a unique IPM $\mu_{\infty}\in\hat{\mathscr{P}}_{2,M}(\mathbb{R}^{d})$. So, the log-Sobolev inequality (2.29) follows from (2.21).
(2) Next, we prove the exponential ergodicity in $\mathbb{W}_{2}$. Let $X_{t}^{i},i=1,2$ solve
$\ \text{\rm{d}}X_{t}^{1}=b(X_{t}^{1},P_{t}^{\ast}\mu_{0})\text{\rm{d}}t+\sigma(P_{t}^{\ast}\mu_{0})\text{\rm{d}}W_{t},$
and
$\text{\rm{d}}X_{t}^{2}=b(X_{t}^{2},\mu_{\infty})\text{\rm{d}}t+\sigma(\mu_{\infty})\text{\rm{d}}W_{t}$
with $X_{0}^{1}=X_{0}^{2}$ satisfying $\mathscr{L}_{X_{0}^{1}}=\mu_{0}$. Similar to (2.34), Itô’s formula and ((H)) imply that
$\mathbb{W}_{2}(P_{t}^{\ast}\mu_{0},(\tilde{P}_{t}^{\mu_{\infty}})^{\ast}\mu_{0})^{2}$ $\leq\mathbb{E}|X_{t}^{1}-X_{t}^{2}|^{2}$
$\leq\int_{0}^{t}\text{\rm{e}}^{K_{1}(t-s)}K_{I}\mathbb{W}_{2}(P_{s}^{\ast}\mu_{0},\mu_{\infty})^{2}\text{\rm{d}}s\leq\text{\rm{e}}^{K_{1}t}K_{I}\int_{0}^{t}\mathbb{W}_{2}(P_{s}^{\ast}\mu_{0},\mu_{\infty})^{2}\text{\rm{d}}s.$
This, together with (2.33) for $\mu=\mu_{\infty},\nu=\mu_{0}$, yields that
$\mathbb{W}_{2}(P_{t}^{\ast}\mu_{0},\mu_{\infty})^{2}$ $\leq 2\mathbb{W}_{2}(P_{t}^{\ast}\mu_{0},(\tilde{P}_{t}^{\mu_{\infty}})^{\ast}\mu_{0})^{2}+2\mathbb{W}_{2}((\tilde{P}_{t}^{\mu_{\infty}})^{\ast}\mu_{0},\mu_{\infty})^{2}$
$\leq 2\text{\rm{e}}^{K_{1}t}K_{I}\int_{0}^{t}\mathbb{W}_{2}(P_{s}^{\ast}\mu_{0},\mu_{\infty})^{2}\text{\rm{d}}s+2c_{1}^{2}\text{\rm{e}}^{-2\lambda_{1}t}\mathbb{W}_{2}(\mu_{0},\mu_{\infty})^{2}.$
Gronwall’s inequality yields
$\mathbb{W}_{2}(P_{t}^{\ast}\mu_{0},\mu_{\infty})^{2}$
(2.35) $\leq\left(\int_{0}^{t}2c_{1}^{2}\text{\rm{e}}^{-2\lambda_{1}s}2\text{\rm{e}}^{K_{1}s}K_{I}\text{\rm{e}}^{2\text{\rm{e}}^{K_{1}t}K_{I}(t-s)}\text{\rm{d}}s+2c_{1}^{2}\text{\rm{e}}^{-2\lambda_{1}t}\right)\mathbb{W}_{2}(\mu_{0},\mu_{\infty})^{2}.$
Take
$\delta_{1}=\sup\left\{K_{I}>0:\inf_{t>\frac{\log(2c_{1}^{2})}{2\lambda_{1}}}\int_{0}^{t}2c_{1}^{2}\text{\rm{e}}^{-2\lambda_{1}s}2\text{\rm{e}}^{K_{1}s}K_{I}\text{\rm{e}}^{2\text{\rm{e}}^{K_{1}t}K_{I}(t-s)}\text{\rm{d}}s+2c_{1}^{2}\text{\rm{e}}^{-2\lambda_{1}t}<1\right\}.$
So, when $K_{I}<\min(K_{2},\delta_{0},\delta_{1})$, we can choose $\hat{t}>0$ and $\alpha\in(0,1)$ such that
(2.36) $\mathbb{W}_{2}(P_{\hat{t}}^{\ast}\mu_{0},\mu_{\infty})\leq\alpha\mathbb{W}_{2}(\mu_{0},\mu_{\infty}).$
By the semigroup property $P_{t+s}^{\ast}=P_{t}^{\ast}P_{s}^{\ast}$ and (2.3)-(2.36), we get (2.30).
(3) Finally, assuming $\sigma(x,\gamma)=\sigma$ and (2.7), the log-Harnack inequality holds by [36], i.e.
$(P_{t}^{\ast}\mu_{0})(\log f)$
(2.37) $\leq\log(P_{t}^{\ast}\nu_{0})(f)+\eta(t)\mathbb{W}_{2}(\mu_{0},\nu_{0})^{2},\ \ t>0,\mu_{0},\nu_{0}\in\mathscr{P}_{2}(\mathbb{R}^{d}),1<f\in\mathscr{B}_{b}(\mathbb{R}^{d})$
holds for some measurable function $\eta:(0,\infty)\rightarrow(0,\infty)$. Moreover, the Talagrand inequality also holds due to the log-Sobolev inequality (2.29), see for instance [27], which together with (2.3) and (2.30) implies (2.32), see [30] for more details. ∎
Remark 2.8.
Compared with the assumption $\sup_{\mu\in\mathscr{P}_{2}(\mathbb{R}^{d})}|b(0,\mu)|<\infty$ in [45, Corollary 2.14], $|b(0,\mu)|$ can be of linear growth in $\mu$ in Theorem 2.7, see Example 2.9 below.
Example 2.9.
Let $V(x)=-|x|^{4}+|x|^{2},x\in\mathbb{R}^{d}$, $W\in C^{2}(\mathbb{R}^{d};\mathbb{R})$, $U\in C^{2}(\mathbb{R}^{d};\mathbb{R}^{d})$ with $\nabla U$ being bounded and there exists a constant $K_{I}>0$ such that
$|\nabla W(x)-\nabla W(y)|+|\nabla U(x)-\nabla U(y)|\leq K_{I}|x-y|,\ \ x,y\in\mathbb{R}^{d}.$
Set
$b(x,\mu)=\nabla V(x)+\int_{\mathbb{R}^{d}}\nabla W(x-y)\mu(\text{\rm{d}}y)~{}\mbox{ and }~{}\sigma(\mu)=\sqrt{I_{d\times d}+\int_{\mathbb{R}^{d}}(\nabla U(\nabla U)^{\ast})(y)\mu(\text{\rm{d}}y)}.$
Then (H) holds and if $K_{I}$ is small enough, the assertions in Theorem 2.7 hold.
Remark 2.10.
If $\sigma$ also depends on the spatial variable and the Bakry-Émery curvature condition holds, i.e.
(2.38) $\Gamma_{2}^{\mu}(f)\geq K|\sigma^{\ast}\nabla f|^{2},\ \ f\in C_{0}^{\infty}(\mathbb{R}^{d}),\mu\in\mathscr{P}_{2}(\mathbb{R}^{d})$
for some constant $K\in\mathbb{R}$, where
$L^{\mu}=\langle b(\cdot,\mu),\nabla\rangle+\frac{1}{2}\mathrm{tr}((\sigma\sigma^{\ast})(\cdot,\mu)\nabla^{2}f),$
and
$\Gamma_{2}^{\mu}(f):=\frac{1}{2}L^{\mu}(|\sigma^{\ast}\nabla f|^{2})-\langle\sigma^{\ast}\nabla f,\sigma^{\ast}\nabla L^{\mu}f\rangle,$
then (2.2) is true, see for instance [2, 15]. Hence, the assertions in Theorem 2.6 and Theorem 2.7 also hold by replacing $\sigma(x,\gamma)=\sigma(\gamma)$ with (2.38). However, (2.38) is not easy to check.
3 Degenerate case
In this section, consider
(3.3) $\left\{\begin{array}[]{ll}\text{\rm{d}}X_{t}=Y_{t}\text{\rm{d}}t,\\ \text{\rm{d}}Y_{t}=b_{t}(X_{t},Y_{t},\mathscr{L}_{(X_{t},Y_{t})})\text{\rm{d}}t+\sigma_{t}(\mathscr{L}_{(X_{t},Y_{t})})\text{\rm{d}}W_{t},\end{array}\right.$
where $b:(0,\infty)\times\mathbb{R}^{2d}\times\mathscr{P}(\mathbb{R}^{2d})\rightarrow\mathbb{R}^{d}$, $\sigma:(0,\infty)\times\mathscr{P}(\mathbb{R}^{2d})\rightarrow\mathbb{R}^{d}\otimes\mathbb{R}^{n}$ are measurable and bounded on bounded sets and $\{W_{t}\}_{t\geq 0}$ is an $n$-dimensional standard Brownian motion on some complete filtered probability space $(\Omega,\mathscr{F},(\mathscr{F}_{t})_{t\geq 0},\mathbb{P})$. As said above, when (3.3) is well-posed in $\mathscr{P}_{2}(\mathbb{R}^{2d})$, we use $P_{t}^{\ast}\mu_{0}$ to denote the distribution of the solution to (3.3) with initial distribution $\mu_{0}\in\mathscr{P}_{2}(\mathbb{R}^{2d})$. For any $\mu_{\cdot}\in C([0,\infty),\mathscr{P}_{2}(\mathbb{R}^{2d}))$, we also consider the time-inhomogeneous decoupled SDEs:
(3.6) $\left\{\begin{array}[]{ll}\text{\rm{d}}X_{t}^{\mu}=Y_{t}^{\mu}\text{\rm{d}}t,\\ \text{\rm{d}}Y_{t}^{\mu}=b_{t}(X_{t}^{\mu},Y_{t}^{\mu},\mu_{t})\text{\rm{d}}t+\sigma_{t}(\mu_{t})\text{\rm{d}}W_{t}.\end{array}\right.$
We make the following assumptions.
  1. (B)
    There exist constants $K_{M}>0,K_{I}>0,K>0$ such that
    $|b_{t}(z,\gamma)-b_{t}(\bar{z},\tilde{\gamma})|+\|\sigma_{t}(\gamma)-\sigma_{t}(\tilde{\gamma})\|_{HS}$
    $\leq K_{M}|z-\bar{z}|+K_{I}\mathbb{W}_{2}(\gamma,\tilde{\gamma}),\ \ z,\bar{z}\in\mathbb{R}^{2d},\gamma,\tilde{\gamma}\in\mathscr{P}_{2}(\mathbb{R}^{2d})$
    and $|b_{t}(0,\delta_{0})|+\|\sigma_{t}(\delta_{0})\|_{HS}\leq K,\ \ t\geq 0.$
Under (B), (3.3) is well-posed in $\mathscr{P}_{2}(\mathbb{R}^{2d})$ and (3.6) is also well-posed. Let $P_{s,t}^{\mu}$ be the semigroup associated to (3.6). For a differentiable function $f$ on $\mathbb{R}^{d}\times\mathbb{R}^{d}$, $i=1,2$, let $\nabla^{(i)}f(x_{1},x_{2})$ denote the gradient of $f$ along $x_{i}$.
3.1 Log-Sobolev inequality for $P_{s,t}^{\mu}$ and $P_{t}^{\ast}\mu_{0}$
Theorem 3.1.
Assume (B) and there exists a constant $\delta_{1}>0$ such that
(3.7) $\sigma_{t}(\gamma)\sigma_{t}(\gamma)^{\ast}\leq\delta_{1},\ \ \gamma\in\mathscr{P}_{2}(\mathbb{R}^{2d}).$
Then the following assertions hold.
  1. (1)
    For any $\mu_{\cdot}\in C([0,\infty),\mathscr{P}_{2}(\mathbb{R}^{2d}))$, the log-Sobolev ineqiality holds: for $0\leq s\leq t,f>0,f\in C_{0}^{\infty}(\mathbb{R}^{2d})$
    (3.8) $P_{s,t}^{\mu}(f\log f)-P_{s,t}^{\mu}(f)\log P_{s,t}^{\mu}(f)\leq\frac{2\delta_{1}}{2K_{M}+1}(\text{\rm{e}}^{(2K_{M}+1)(t-s)}-1)P_{s,t}^{\mu}(|\nabla f^{\frac{1}{2}}|^{2}).$
  2. (2)
    If $\mu_{0}$ satisfies the log-Sobolev inequality: for some $L_{0}>0$,
    $\mu_{0}(f\log f)\leq\mu_{0}(f)\log\mu_{0}(f)+2L_{0}\mu_{0}(|\nabla f^{\frac{1}{2}}|^{2}),\ \ f>0,f\in C_{0}^{\infty}(\mathbb{R}^{2d})$
    then $P_{t}^{\ast}\mu_{0}$ also admits the log-Sobolev inequality
    $(P_{t}^{\ast}\mu_{0})(f\log f)-(P_{t}^{\ast}\mu_{0})(f)\log(P_{t}^{\ast}\mu_{0})(f)$
    $\leq 2\left\{\frac{\delta_{1}}{2K_{M}+1}(\text{\rm{e}}^{(2K_{M}+1)t}-1)+L_{0}\text{\rm{e}}^{(2K_{M}+1)t}\right\}(P_{t}^{\ast}\mu_{0})(|\nabla f^{\frac{1}{2}}|^{2}),$
    $\ \ f>0,f\in C_{0}^{\infty}(\mathbb{R}^{2d}),t\geq 0.$
Proof.
Let $\{X_{s,t}^{\mu,z}\}_{0\leq s\leq t}$ be the solution to (3.6) with $X_{s,s}^{\mu,z}=z\in\mathbb{R}^{2d}$. By (B), it is standard to derive that
$|X_{s,t}^{\mu,z}-X_{s,t}^{\mu,\bar{z}}|\leq\text{\rm{e}}^{(K_{M}+\frac{1}{2})(t-s)}|z-\bar{z}|,\ \ z,\bar{z}\in\mathbb{R}^{2d},0\leq s\leq t.$
This implies that
(3.9) $|\nabla P_{s,t}^{\mu}f|\leq\text{\rm{e}}^{(K_{M}+\frac{1}{2})(t-s)}P_{s,t}^{\mu}|\nabla f|,\ \ f\in C_{0}^{\infty}(\mathbb{R}^{2d}),0\leq s\leq t.$
In the present case, we have $\Gamma^{t,\mu}_{1}(f)=|(\sigma_{t}^{\mu})^{\ast}\nabla^{(2)}f|^{2}$ and it is easy to see that
$\Gamma_{1}^{r,\mu}(P_{r,t}^{\mu}f)$ $=|(\sigma_{r}^{\mu})^{\ast}\nabla^{(2)}P_{r,t}^{\mu}f|^{2}\leq\delta_{1}|\nabla P_{r,t}^{\mu}f|^{2},\ \ f\in C_{0}^{\infty}(\mathbb{R}^{2d}),0\leq r\leq t.$
The remaining is completely the same with that of Theorem 2.1 with $K_{1}$ replaced by $12K_{M}+$ 1 and we omit it. ∎
Remark 3.2.
It is not difficult to see that (3.8) does not hold if $|\nabla f|^{2}$ is replaced by $|\nabla^{(2)}f|^{2}$. This is essentially different from the non-degenerate case. In fact, in the proof of Theorem 3.1, (3.9) does not hold if $\nabla$ is replaced by $\nabla^{(2)}$.
3.2 Hyperboundedness for time-homogenenous decoupled SDEs
From now on, we consider the time-homogeneous case, i.e. consider
(3.12) $\left\{\begin{array}[]{ll}\text{\rm{d}}X_{t}=Y_{t}\text{\rm{d}}t,\\ \text{\rm{d}}Y_{t}=b(X_{t},Y_{t},\mathscr{L}_{(X_{t},Y_{t})})\text{\rm{d}}t+\sigma(\mathscr{L}_{(X_{t},Y_{t})})\text{\rm{d}}W_{t}.\end{array}\right.$
We present the following condition.
  1. (C)
    There exist constants $K_{M}>0,K_{I}>0$ such that for all $z,\bar{z}\in\mathbb{R}^{2d},\gamma,\tilde{\gamma}\in\mathscr{P}_{2}(\mathbb{R}^{2d})$,
    (3.13) $|b(z,\gamma)-b(\bar{z},\tilde{\gamma})|+\|\sigma(\gamma)-\sigma(\tilde{\gamma})\|_{HS}\leq K_{M}|z-\bar{z}|+K_{I}\mathbb{W}_{2}(\gamma,\tilde{\gamma}).$
    Moreover, there exist constants $\theta,r,R\geq 0,$ and $r_{0}\in(-1,1)$ such that for any $x,y,\bar{x},\bar{y}\in\mathbb{R}^{d}$ satisfying $|x-\bar{x}|^{2}+|y-\bar{y}|^{2}\geq R^{2}$ and $\mu\in\mathscr{P}_{2}(\mathbb{R}^{2d})$,
    $\langle r^{2}(x-\bar{x})+rr_{0}(y-\bar{y}),y-\bar{y}\rangle$
    (3.14) $+\langle(y-\bar{y})+rr_{0}(x-\bar{x}),b(x,y,\mu)-b(\bar{x},\bar{y},\mu)\rangle$
    $\leq-\theta(|x-\bar{x}|^{2}+|y-\bar{y}|^{2}).$
((C)) with $R=0$ means uniform dissipativity while ((C)) with $R>0$ becomes partial dissipativity. Let
$\psi((x,y),(\bar{x},\bar{y}))=\sqrt{r^{2}|x-\bar{x}|^{2}/2+|y-\bar{y}|^{2}/2+rr_{0}\langle x-\bar{x},y-\bar{y}\rangle},\ \ (x,y),(\bar{x},\bar{y})\in\mathbb{R}^{2d}.$
Then there exists a constant $C_{\psi}>1$ depending on $r,r_{0}$ such that for any $(x,y),(\bar{x},\bar{y})\in\mathbb{R}^{2d}$,
(3.15) $C_{\psi}^{-1}(|x-\bar{x}|^{2}+|y-\bar{y}|^{2})\leq\psi((x,y),(\bar{x},\bar{y}))^{2}\leq C_{\psi}(|x-\bar{x}|^{2}+|y-\bar{y}|^{2}).$
For any $\mu\in\mathscr{P}_{2}(\mathbb{R}^{2d})$, we first study time-homogeneous decoupled SDEs
(3.18) $\left\{\begin{array}[]{ll}\text{\rm{d}}\tilde{X}_{t}^{\mu}=\tilde{Y}_{t}^{\mu}\text{\rm{d}}t,\\ \text{\rm{d}}\tilde{Y}_{t}^{\mu}=b(\tilde{X}_{t}^{\mu},\tilde{Y}_{t}^{\mu},\mu)\text{\rm{d}}t+\sigma(\mu)\text{\rm{d}}W_{t}.\end{array}\right.$
Let $\tilde{P}_{t}^{\mu}$ be the associated semigroup to (3.18) and $(\tilde{P}_{t}^{\mu})^{\ast}\gamma$ stand for the distribution of the solution to (3.18) with initial distribution $\gamma\in\mathscr{P}_{2}(\mathbb{R}^{2d})$. To derive the hyperboundedness of $\tilde{P}_{t}^{\mu}$, we first investigate the concentration of the invariant probability measure to (3.18).
Theorem 3.3.
Assume (C) and there is a constat $\delta_{2}>0$ such that
(3.19) $\sigma(\gamma)\sigma^{\ast}(\gamma)\geq\delta_{2},\ \ \gamma\in\mathscr{P}_{2}(\mathbb{R}^{2d}).$
Then for any $\mu\in\mathscr{P}_{2}(\mathbb{R}^{2d})$, (3.18) has a unique invariant probability measure denoted by $\Phi(\mu)$. Moreover, there exist constants $N>0$ and $\eta_{0}>0$ such that when $K_{I}<\eta_{0}$, $\Phi$ is a map on $\hat{\mathscr{P}}_{2,N}(\mathbb{R}^{2d})$ with
$\hat{\mathscr{P}}_{2,N}(\mathbb{R}^{2d}):=\left\{\mu\in\mathscr{P}_{2}(\mathbb{R}^{2d}):\|\mu\|_{2}^{2}\leq N\right\}.$
In addition, there exist constants $\varepsilon>0,C_{0}>0$ such that
(3.20) $\{\Phi(\mu)\}(\text{\rm{e}}^{\varepsilon|\cdot|^{2}})<C_{0},\ \ \mu\in\hat{\mathscr{P}}_{2,N}(\mathbb{R}^{2d}).$
Proof.
(1) Firstly, it follows from (3.13) that
(3.21) $|b(0,\mu)|+\|\sigma(\mu)\|_{HS}\leq K_{I}\|\mu\|_{2}+|b(0,\delta_{0})|+|\sigma(\delta_{0})|,\ \ \mu\in\mathscr{P}_{2}(\mathbb{R}^{2d}).$
Let
$\tilde{\psi}(x,y)=\psi((x,y),0)=\sqrt{r^{2}|x|^{2}/2+|y|^{2}/2+rr_{0}\langle x,y\rangle},\ \ (x,y)\in\mathbb{R}^{2d}.$
It follows from Itô’s formula, ((C)) and (3.21) that for some positive constants $c_{1},c_{2},c_{3}$,
$\text{\rm{d}}\tilde{\psi}(\tilde{X}_{t}^{\mu},\tilde{Y}_{t}^{\mu})^{2}$ $=\langle r^{2}\tilde{X}_{t}^{\mu}+rr_{0}\tilde{Y}_{t}^{\mu},\tilde{Y}_{t}^{\mu}\rangle\text{\rm{d}}t+\langle\tilde{Y}_{t}^{\mu}+rr_{0}\tilde{X}_{t}^{\mu},b(\tilde{X}_{t}^{\mu},\tilde{Y}_{t}^{\mu},\mu)\rangle\text{\rm{d}}t$
$\qquad\quad+\frac{1}{2}\|\sigma(\mu)\|_{HS}^{2}\text{\rm{d}}t+\langle\tilde{Y}_{t}^{\mu}+rr_{0}\tilde{X}_{t}^{\mu},\sigma(\mu)\text{\rm{d}}W_{t}\rangle$
(3.22) $\leq-\theta(|\tilde{X}_{t}^{\mu}|^{2}+|\tilde{Y}_{t}^{\mu}|^{2})\text{\rm{d}}t+\left(\theta+\frac{r^{2}}{2}+|rr_{0}|+K_{M}(1+|rr_{0}|)\right)R^{2}\text{\rm{d}}t$
$\qquad\quad+\langle\tilde{Y}_{t}^{\mu}+rr_{0}\tilde{X}_{t}^{\mu},b(0,\mu)\rangle\text{\rm{d}}t+\frac{1}{2}\|\sigma(\mu)\|_{HS}^{2}\text{\rm{d}}t+\langle\tilde{Y}_{t}^{\mu}+rr_{0}\tilde{X}_{t}^{\mu},\sigma(\mu)\text{\rm{d}}W_{t}\rangle$
$\leq-c_{2}(|\tilde{X}_{t}^{\mu}|^{2}+|\tilde{Y}_{t}^{\mu}|^{2})\text{\rm{d}}t+c_{1}\text{\rm{d}}t+c_{3}K_{I}^{2}\|\mu\|_{2}^{2}\text{\rm{d}}t+\langle\tilde{Y}_{t}^{\mu}+rr_{0}\tilde{X}_{t}^{\mu},\sigma(\mu)\text{\rm{d}}W_{t}\rangle.$
So, we arrive at
(3.23) $\frac{1}{t}\int_{0}^{t}\|(\tilde{P}_{s}^{\mu})^{\ast}\delta_{0}\|_{2}^{2}\text{\rm{d}}s\leq(c_{1}+c_{3}K_{I}^{2}\|\mu\|_{2}^{2})/{c_{2}}.$
By the standard argument of tightness, we derive from (3.23) that (3.18) has an invariant probability measure denoted by $\Phi(\mu)$, which satisfies
(3.24) $\|\Phi(\mu)\|_{2}^{2}\leq(c_{1}+c_{3}K_{I}^{2}\|\mu\|_{2}^{2})/{c_{2}}.$
Moreover, by (3.19), (3.13) and [34, Lemma 3.2], the Harnack inequality for $\tilde{P}_{t}^{\mu}$ with power $2$ holds, i.e. for any $t>0$, there exists a constant $c_{0}>0$ independent of $\mu$ such that
(3.25) $(\tilde{P}_{t}^{\mu}f)^{2}(z)\leq\tilde{P}_{t}^{\mu}f^{2}(\bar{z})\text{\rm{e}}^{c_{0}|z-\bar{z}|^{2}},\ \ f\in\mathscr{B}_{b}^{+}(\mathbb{R}^{2d}),z,\bar{z}\in\mathbb{R}^{2d},\mu\in\mathscr{P}_{2}(\mathbb{R}^{2d}).$
This implies the uniqueness of the invariant probability measure of (3.18), see for instance [35, Theorem 1.4.1].
(2) Next let $K_{I}<\sqrt{\frac{c_{2}}{c_{3}}}$. For any $\mu\in\mathscr{P}_{2}(\mathbb{R}^{2d})$ with $\|\mu\|_{2}^{2}\leq c_{1}(c_{2}-c_{3}K_{I}^{2})^{-1}$, it follows from (3.24) that
$\|\Phi(\mu)\|_{2}^{2}\leq(c_{1}+c_{3}K_{I}^{2}\|\mu\|_{2}^{2})/{c_{2}}\leq c_{1}(c_{2}-c_{3}K_{I}^{2})^{-1}.$
This means that $\Phi$ is a map on $\hat{\mathscr{P}}_{2}(\mathbb{R}^{2d})$ with
$\hat{\mathscr{P}}_{2}(\mathbb{R}^{2d}):=\left\{\mu\in\mathscr{P}_{2}(\mathbb{R}^{2d}):\|\mu\|_{2}^{2}\leq c_{1}(c_{2}-c_{3}K_{I}^{2})^{-1}\right\}.$
(3) Finally, we study the concentration of $\Phi(\mu)$ for any $\mu\in\hat{\mathscr{P}}_{2}(\mathbb{R}^{2d})$. By (3.2), (3.21) and Itô’s formula for $\text{\rm{e}}^{\varepsilon\tilde{\psi}(\tilde{X}_{t}^{\mu},\tilde{Y}_{t}^{\mu})^{2}}$, for any $\mu\in\hat{\mathscr{P}}_{2}(\mathbb{R}^{2d})$, we derive
$\text{\rm{d}}\text{\rm{e}}^{\varepsilon\tilde{\psi}(\tilde{X}_{t}^{\mu},\tilde{Y}_{t}^{\mu})^{2}}$
$\leq\varepsilon\text{\rm{e}}^{\varepsilon\tilde{\psi}(\tilde{X}_{t}^{\mu},\tilde{Y}_{t}^{\mu})^{2}}\left\{-c_{2}(|\tilde{X}_{t}^{\mu}|^{2}+|\tilde{Y}_{t}^{\mu}|^{2})\text{\rm{d}}t+c_{1}\text{\rm{d}}t+c_{3}K_{I}^{2}\|\mu\|_{2}^{2}\text{\rm{d}}t+\langle\tilde{Y}_{t}^{\mu}+rr_{0}\tilde{X}_{t}^{\mu},\sigma(\mu)\text{\rm{d}}W_{t}\rangle\right\}$
$+\frac{1}{2}\varepsilon^{2}\text{\rm{e}}^{\varepsilon\tilde{\psi}(\tilde{X}_{t}^{\mu},\tilde{Y}_{t}^{\mu})^{2}}|\sigma(\mu)^{\ast}[\tilde{Y}_{t}^{\mu}+rr_{0}\tilde{X}_{t}^{\mu}]|^{2}$
$\leq\varepsilon\text{\rm{e}}^{\varepsilon\tilde{\psi}(\tilde{X}_{t}^{\mu},\tilde{Y}_{t}^{\mu})^{2}}\left\{(-c_{2}+\varepsilon c_{4})(|\tilde{X}_{t}^{\mu}|^{2}+|\tilde{Y}_{t}^{\mu}|^{2})\text{\rm{d}}t+c_{5}\text{\rm{d}}t+\langle\tilde{Y}_{t}^{\mu}+rr_{0}\tilde{X}_{t}^{\mu},\sigma(\mu)\text{\rm{d}}W_{t}\rangle\right\}$
for some positive constants $c_{4},c_{5}$. When $\varepsilon>0$ is small enough, we can find constants $c_{6}(\varepsilon),c_{7}(\varepsilon)>0$ such that
$\varepsilon\text{\rm{e}}^{\varepsilon\tilde{\psi}(\tilde{X}_{t}^{\mu},\tilde{Y}_{t}^{\mu})^{2}}\left\{(-c_{2}+\varepsilon c_{4})(|\tilde{X}_{t}^{\mu}|^{2}+|\tilde{Y}_{t}^{\mu}|^{2})+c_{5}\right\}\leq c_{6}(\varepsilon)-c_{7}(\varepsilon)\text{\rm{e}}^{\varepsilon\tilde{\psi}(\tilde{X}_{t}^{\mu},\tilde{Y}_{t}^{\mu})^{2}}.$
This together with (3.15) yields $\frac{1}{t}\int_{0}^{t}\{(\tilde{P}_{s}^{\mu})^{\ast}\delta_{0}\}(\text{\rm{e}}^{\varepsilon C_{\psi}^{-1}|\cdot|^{2}})\text{\rm{d}}s\leq\frac{c_{6}(\varepsilon)t+1}{c_{7}(\varepsilon)t}.$ Thus, the assertion (3.20) is derived by following the line to derive (3.24). ∎
The Harnack inequality (3.25) is not as sharp as the one in (2.2). In the following theorem, we will first derive a sharper Harnack inequality, which implies the hyperboundedness of $\tilde{P}_{t}^{\mu}$ for large enough $t$.
Theorem 3.4.
Assume (C) and (3.19). Then
$(\tilde{P}_{t}^{\mu}f)^{2}(z)\leq\tilde{P}_{t}^{\mu}f^{2}(\bar{z})$
(3.26) ×$\quad\times\text{\rm{e}}^{c_{0}C_{1}\frac{C_{\psi}^{2}}{\theta}+c_{0}C_{\psi}^{2}\text{\rm{e}}^{-\frac{\theta}{C_{\psi}}(t-1)}|z-\bar{z}|^{2}},\ \ t\geq 1,f\in\mathscr{B}_{b}^{+}(\mathbb{R}^{2d}),z,\bar{z}\in\mathbb{R}^{2d},\mu\in\mathscr{P}_{2}(\mathbb{R}^{2d})$
holds for some constant $C_{1}>0$ depending on $r,r_{0},\theta,R,K_{M}$. Consequently, if $K_{I}$ is small enough, then there exist constants $t_{0}>0,\tilde{C}>0$ such that the hyperboundedness of $\tilde{P}_{t_{0}}^{\mu}$ holds, i.e.
(3.27) $\|\tilde{P}_{t_{0}}^{\mu}\|_{L^{2}(\Phi(\mu))\rightarrow L^{4}(\Phi(\mu))}\leq\tilde{C},\ \ \mu\in\hat{\mathscr{P}}_{2,N}(\mathbb{R}^{2d}).$
Proof.
Let $(X_{t},Y_{t})$ and $(\bar{X}_{t},\bar{Y}_{t})$ solve (3.18) with $(X_{0},Y_{0})=z,(\bar{X}_{0},\bar{Y}_{0})=\bar{z}$. By Itô’s formula, (3.13), ((C)) and (3.15), we can find a constant $C_{1}>0$ depending on $r,r_{0},\theta,R,K_{M}$ such that
$\text{\rm{d}}\psi((X_{t},Y_{t}),(\bar{X}_{t},\bar{Y}_{t}))^{2}$ $=\langle r^{2}(X_{t}-\bar{X}_{t})+rr_{0}(Y_{t}-\bar{Y}_{t}),Y_{t}-\bar{Y}_{t}\rangle$
$+\langle(Y_{t}-\bar{Y}_{t})+rr_{0}(X_{t}-\bar{X}_{t}),b(X_{t},Y_{t},\mu)-b(\bar{X}_{t},\bar{Y}_{t},\mu)\rangle$
$\leq-\theta(|X_{t}-\bar{X}_{t}|^{2}+|Y_{t}-\bar{Y}_{t}|^{2})\text{\rm{d}}t+C_{1}\text{\rm{d}}t$
$\leq-\frac{\theta}{C_{\psi}}(|X_{t}-\bar{X}_{t}|^{2}+|Y_{t}-\bar{Y}_{t}|^{2})\text{\rm{d}}t+C_{1}\text{\rm{d}}t.$
This again combined with (3.15) implies that
(3.28) $|X_{t}-\bar{X}_{t}|^{2}+|Y_{t}-\bar{Y}_{t}|^{2}\leq C_{1}\frac{C_{\psi}^{2}}{\theta}+C_{\psi}^{2}\text{\rm{e}}^{-\frac{\theta}{C_{\psi}}t}|z-\bar{z}|^{2}.$
On the other hand, it follows from (3.25) with $t=1$ and (3.28) that
$(\tilde{P}_{1}^{\mu}f)^{2}(X_{t-1})$ $\leq\tilde{P}_{1}^{\mu}f^{2}(\bar{X}_{t-1})\text{\rm{e}}^{c_{0}(|X_{t-1}-\bar{X}_{t-1}|^{2}+|Y_{t-1}-\bar{Y}_{t-1}|^{2})}$
(3.29) $\leq\tilde{P}_{1}^{\mu}f^{2}(\bar{X}_{t-1})\text{\rm{e}}^{c_{0}C_{1}\frac{C_{\psi}^{2}}{\theta}+c_{0}C_{\psi}^{2}\text{\rm{e}}^{-\frac{\theta}{C_{\psi}}(t-1)}|z-\bar{z}|^{2}},\ \ t\geq 1,f\in\mathscr{B}_{b}^{+}(\mathbb{R}^{2d}).$
Taking expectation in (3.2) and using the semigroup property and Jensen’s inequality, we deduce (3.4). Repeating the same argument to derive (2.26) from (2.2) and (2.20), we gain (3.27) from (3.20) and (3.4). ∎
3.3 Log-Sobolev and Talagrand inequality for $\Phi(\mu)$
Theorem 3.5.
Assume (C), (3.7), (3.19) and $K_{I}$ is small enough.
  1. 1.
    Then for any $\mu\in\hat{\mathscr{P}}_{2,N}(\mathbb{R}^{2d})$, $\Phi(\mu)$ satisfies the log-Sobolev inequality
    (3.30) $\Phi(\mu)(f^{2}\log f^{2})\leq C_{L}(\Phi(\mu))\Phi(\mu)(|\nabla f|^{2}),\ \ \Phi(\mu)(f^{2})=1,\ \ f\in C_{0}^{\infty}(\mathbb{R}^{2d})$
    for some constant $C_{L}(\Phi(\mu))>0$ depending on $\Phi(\mu)$. Consequently, for any $\mu\in\hat{\mathscr{P}}_{2,N}(\mathbb{R}^{2d})$, the Talagrand inequality
    (3.31) $\mathbb{W}_{2}(\mu_{0},\Phi(\mu))^{2}\leq\tilde{C}_{L}(\Phi(\mu))\mathrm{Ent}(\mu_{0}|\Phi(\mu)),\ \ \mu_{0}\in\mathscr{P}_{2}(\mathbb{R}^{2d}).$
    holds for some constant $\tilde{C}_{L}(\Phi(\mu))>0$.
  2. 2.
    If moreover, for any $R>0$ and $N>0$, there exist positive constants $C_{1}$ and $C_{2}$ depending on $R$, $N$ such that
    (3.32) $C_{1}\leq\rho_{\mu}(x)\leq C_{2},\ \ x\in B_{R},\mu\in\hat{\mathscr{P}}_{2,N}(\mathbb{R}^{2d})$
    for $\rho_{\mu}:=\frac{\text{\rm{d}}\Phi(\mu)}{\text{\rm{d}}x}$. Then there exists a constant $\bar{C}>0$ depending on $N$ such that
    $\Phi(\mu)(f^{2}\log f^{2})\leq\bar{C}\Phi(\mu)(|\nabla f|^{2}),$
    (3.33) $\qquad\quad\Phi(\mu)(f^{2})=1,\ \ f\in C_{0}^{\infty}(\mathbb{R}^{2d}),\mu\in\hat{\mathscr{P}}_{2,N}(\mathbb{R}^{2d}),$
    and
    (3.34) $\mathbb{W}_{2}(\mu_{0},\Phi(\mu))^{2}\leq\bar{C}\mathrm{Ent}(\mu_{0}|\Phi(\mu)),\ \ \mu_{0}\in\mathscr{P}_{2}(\mathbb{R}^{2d}),\mu\in\hat{\mathscr{P}}_{2,N}(\mathbb{R}^{2d}).$
Proof.
Combining the hyperboundedness (3.27) and the log-Sobolev inequality (3.8) with $s=0$ and $\tilde{P}_{t}^{\mu}$ replacing $P_{s,t}^{\mu}$, we obtain a defective log-Sobolev inequality by Lemma 2.5(1), i.e. there exist constant $\tilde{C}_{1},\tilde{C}_{2}>0$ independent of $\mu$ such that
$\Phi(\mu)(f\log f)-\Phi(\mu)(f)\log\Phi(\mu)(f)$
[…]
The irreducibility of $\mathscr{E}(f,f)=\Phi(\mu)(|\nabla f|^{2})$ implies (3.30) by [37, Corollary 1.3]. This further implies (3.31) according to [27]. Finally, by (3.3), (3.20) and (3.32), the proof is completed by Lemma 2.5(2). ∎
Remark 3.6.
The drift $b$ in Theorem 3.5 is allowed to be partially dissipative, which improves the existing results in [30] where $b$ is assumed to be uniformly dissipative. In contrast to the results in [32] and [4], the drift is allowed to be not of gradient-type. We should also point out that different from the non-degenerate case, the log-Sobolev inequality (3.30) does not imply the exponential convergence of $\tilde{P}_{t}^{\mu}$ in relative entropy since the Dirichlet form is $\Phi(\mu)(|\nabla f|^{2})$ instead of $\Phi(\mu)(|\nabla^{(2)}f|^{2})$.
3.4 Exponential ergodicity of (3.12)
Finally, we will use Theorem 3.5 to study the exponential ergodicity of (3.12). We divide it into two cases: uniformly dissipative case (((C)) with $R=0$) and partially dissipative case(((C)) with $R>0$). In the former one, $b$ can be general while in the latter one, $b$ is assumed to be of gradient-type.
3.4.1 Uniformly dissipative case
Theorem 3.7.
Assume (C) with $R=0$, (3.7) and (3.19). If $K_{I}$ is small enough, then (3.12) has a unique IPM $\bar{\mu}\in\mathscr{P}_{2}(\mathbb{R}^{2d})$, and there exist constants $c,\lambda>0$ such that
(3.36) $\mathbb{W}_{2}(P_{t}^{\ast}\mu_{0},\bar{\mu})\leq c\text{\rm{e}}^{-\lambda t}\mathbb{W}_{2}(\mu_{0},\bar{\mu}),\ \ t\geq 0,\mu_{0}\in\mathscr{P}_{2}(\mathbb{R}^{2d}).$
Moreover, there exists a constant $\tilde{c}>0$ such that
(3.37) $\mathrm{Ent}(P_{t}^{\ast}\mu_{0}|\bar{\mu})\leq\tilde{c}^{2}\text{\rm{e}}^{-2\lambda t}\mathrm{Ent}(\mu_{0}|\bar{\mu}),~{}t\geq 1,\mu_{0}\in\mathscr{P}_{2}(\mathbb{R}^{2d}).$
Remark 3.8.
(3.37) in Theorem 3.7 extends the results in [30, Theorem 2.4] and [17, Theorem 4.2], where $b$ is assumed to be of gradient-type, i.e. $b(x,y,\mu)=-y-x-\nabla V(x,\mu)$.
Proof.
By (C) with $R=0$, it follows from [17, Theorem 4.2] that (3.12) has a unique IPM $\bar{\mu}\in\mathscr{P}_{2}(\mathbb{R}^{d})$ satisfying (3.36). Since $\Phi(\bar{\mu})=\bar{\mu}\in\hat{\mathscr{P}}_{2,N}(\mathbb{R}^{2d})$, (3.31) implies the following Talagrand inequality
(3.38) $\mathbb{W}_{2}(\mu_{0},\bar{\mu})^{2}\leq\tilde{C}_{L}\mathrm{Ent}(\mu_{0}|\bar{\mu}),\ \ \mu_{0}\in\mathscr{P}_{2}(\mathbb{R}^{2d}).$
By [17, Theorem 3.1], there exists a measurable function $c\colon(0,\infty)\rightarrow(0,\infty)$ such that
$(P_{t}^{\ast}\mu_{0})(\log f)$
(3.39) $\leq\log(P_{t}^{\ast}\nu_{0})(f)+c(t)\mathbb{W}_{2}(\mu_{0},\nu_{0})^{2},\ \ t>0,\mu_{0},\nu_{0}\in\mathscr{P}_{2}(\mathbb{R}^{2d}),0<f\in\mathscr{B}_{b}(\mathbb{R}^{2d}).$
Combining (3.4.1), (3.36) and (3.38), we derive (3.37) by the same argument we derived (2.32). ∎
3.4.2 Partially dissipative case
Theorem 3.9.
Assume ((C)) with $R>0$, $\sigma=I_{d\times d}$ and
$b(x,y,\mu)=-y-\nabla V(x)-\int_{\mathbb{R}^{2d}}\nabla_{x}W(x,z)\mu(\text{\rm{d}}z),\ \ x,y\in\mathbb{R}^{d}$
with $V:\mathbb{R}^{d}\rightarrow\mathbb{R}$ and $W:\mathbb{R}^{d}\times\mathbb{R}^{2d}\rightarrow\mathbb{R}$ satisfying
(3.40) $\|\nabla^{2}V\|+\|\nabla_{x}^{2}W\|\leq K_{M}\quad\mbox{ and }\quad\|\nabla_{z}\nabla_{x}W\|\leq K_{I}$
for some constants $K_{M},K_{I}>0$. If $K_{I}$ is small enough, then (3.12) has a unique invariant probability measure $\bar{\mu}\in\mathscr{P}_{2}(\mathbb{R}^{2d})$. Moreover, there exist constants $c>0,\lambda>0$ such that
(3.41) $\mathbb{W}_{2}(P_{t}^{\ast}\mu_{0},\bar{\mu})\leq c\text{\rm{e}}^{-\lambda t}\mathbb{W}_{2}(\mu_{0},\bar{\mu}),\ \ t\geq 0,$
and
$\mathrm{Ent}(P_{t}^{\ast}\mu_{0}|\bar{\mu})\leq c^{2}\text{\rm{e}}^{-2\lambda t}\mathrm{Ent}(\mu_{0}|\bar{\mu}),\ \ t\geq 1.$
Remark 3.10.
In the distribution independent case, [32, Theorem 39] derives the exponential ergodicity in relative entropy by means of the log-Sobolev inequality for the explicit invariant probability measure on the product space $\mathbb{R}^{2d}$. In Theorem 3.9, we offer an explicit partially dissipative condition under which the log-Sobolev inequality holds. Moreover, the authors in [14] adopt the log-Sobolev inequality and log-Harnack inequality for mean field interacting particle system as well as propagation of chaos to deduce the exponential ergodicity in $L^{2}$-Wasseretein distance and mean field entropy.
Proof.
(3.12) and (3.18) reduce to
(3.44) $\left\{\begin{array}[]{ll}\text{\rm{d}}X_{t}=Y_{t}\text{\rm{d}}t,\\ \text{\rm{d}}Y_{t}=-Y_{t}\text{\rm{d}}t-\nabla V(X_{t})\text{\rm{d}}t-\int_{\mathbb{R}^{2d}}\nabla_{x}W(X_{t},z)\mathscr{L}_{(X_{t},Y_{t})}(\text{\rm{d}}z)\text{\rm{d}}t+\text{\rm{d}}W_{t},\end{array}\right.$
and
$\left\{\begin{array}[]{ll}\text{\rm{d}}\tilde{X}_{t}^{\mu}=\tilde{Y}_{t}^{\mu}\text{\rm{d}}t,\\ \text{\rm{d}}\tilde{Y}_{t}^{\mu}=-\tilde{Y}_{t}^{\mu}\text{\rm{d}}t-\nabla V(\tilde{X}_{t}^{\mu})\text{\rm{d}}t-\int_{\mathbb{R}^{2d}}\nabla_{x}W(\tilde{X}_{t}^{\mu},z)\mu(\text{\rm{d}}z)\text{\rm{d}}t+\text{\rm{d}}W_{t}.\end{array}\right.$
Note that (3.40) implies (3.13), which together with ((C)) means (C) holds. Observe that the invariant probability measure is
$\rho_{\mu}(x,y):=\frac{\text{\rm{d}}\Phi(\mu)}{\text{\rm{d}}x\text{\rm{d}}y}=Z_{0}\exp\left\{-\frac{1}{2}|y|^{2}-V(x)-\int_{\mathbb{R}^{2d}}W(x,z)\mu(\text{\rm{d}}z)\right\}\text{\rm{d}}x\text{\rm{d}}y,$
where
$Z_{0}=\int_{\mathbb{R}^{2d}}\exp\left\{-\frac{1}{2}|y|^{2}-V(x)-\int_{\mathbb{R}^{2d}}W(x,z)\mu(\text{\rm{d}}z)\right\}\text{\rm{d}}x\text{\rm{d}}y.$
Since $V$ and $W$ are Lipschitz continuous, we conclude that for any $R>0$ and $N>0$, there exist constant $C_{1}$ and $C_{2}$ depending on $R$, $N$ such that
$C_{1}\leq\rho_{\mu}(x,y)\leq C_{2},\ \ (x,y)\in B_{R},\mu\in\hat{\mathscr{P}}_{2,N}(\mathbb{R}^{2d}).$
So, by Theorem 3.5, when $K_{I}$ is small enough, we obtain (2), which combined with [32, Theorem 39] implies that we can find constants $c,\lambda>0$ such that
(3.45) $\mathrm{Ent}((\tilde{P}_{t}^{\mu})^{\ast}\mu_{0}|\Phi(\mu))\leq c\text{\rm{e}}^{-2\lambda t}\mathrm{Ent}(\mu_{0}|\Phi(\mu)),\ \ t\geq 1,\mu_{0}\in\mathscr{P}_{2}(\mathbb{R}^{2d}),\mu\in\hat{\mathscr{P}}_{2,N}(\mathbb{R}^{2d}).$
Note that (3.40) yields that for any $t>0$, there exists a constant $c_{0}>0$ such that
$\tilde{P}_{t}^{\mu}\log f(z)\leq\log\tilde{P}_{t}^{\mu}f(\bar{z})+c_{0}|z-\bar{z}|^{2},\ \ z,\bar{z}\in\mathbb{R}^{2d},f>0,f\in\mathscr{B}_{b}(\mathbb{R}^{2d}),\mu\in\mathscr{P}_{2}(\mathbb{R}^{2d}),$
see for instance [30]. Combining this with (3.34) and (3.45), we conclude
$\mathbb{W}_{2}((\tilde{P}_{t}^{\mu})^{\ast}\mu_{0},\Phi(\mu))^{2}$ $\leq\tilde{C}_{L}\mathrm{Ent}((\tilde{P}_{t-1}^{\mu})^{\ast}[(\tilde{P}_{1}^{\mu})^{\ast}\mu_{0}]|\Phi(\mu))$
$\leq\tilde{c}_{0}\text{\rm{e}}^{-2\lambda t}\mathrm{Ent}((\tilde{P}_{1}^{\mu})^{\ast}\mu_{0}|\Phi(\mu))$
$\leq\tilde{c}_{1}\text{\rm{e}}^{-2\lambda t}\mathbb{W}_{2}(\mu_{0},\Phi(\mu))^{2},\ \ t\geq 2,\mu_{0}\in\mathscr{P}_{2}(\mathbb{R}^{2d}),\mu\in\hat{\mathscr{P}}_{2,N}(\mathbb{R}^{2d}).$
By (3.40), it is easy to find a constant $c_{2}>0$ such that
$\mathbb{W}_{2}((\tilde{P}_{t}^{\mu})^{\ast}\mu_{0},\Phi(\mu))^{2}\leq c_{2}\mathbb{W}_{2}(\mu_{0},\Phi(\mu))^{2},\ \ t\in[0,2],\mu_{0}\in\mathscr{P}_{2}(\mathbb{R}^{2d}),\mu\in\mathscr{P}_{2}(\mathbb{R}^{2d}).$
So, we have
$\mathbb{W}_{2}((\tilde{P}_{t}^{\mu})^{\ast}\mu_{0},\Phi(\mu))^{2}$
(3.46) $\leq\tilde{c}_{2}\text{\rm{e}}^{-2\lambda t}\mathbb{W}_{2}(\mu_{0},\Phi(\mu))^{2},\ \ t\geq 0,\mu_{0}\in\mathscr{P}_{2}(\mathbb{R}^{2d}),\mu\in\hat{\mathscr{P}}_{2,N}(\mathbb{R}^{2d}).$
Again by (3.40) and Gronwall’s inequality, we obtain
(3.47) $\mathbb{W}_{2}((\tilde{P}_{t}^{\mu})^{\ast}\Phi(\nu),(\tilde{P}_{t}^{\nu})^{\ast}\Phi(\nu))^{2}\leq\text{\rm{e}}^{(2+2K_{M})t}\int_{0}^{t}K_{I}^{2}\mathbb{W}_{2}(\mu,\nu)^{2}\text{\rm{d}}s,\ \ \mu,\nu\in\mathscr{P}_{2}(\mathbb{R}^{2d}).$
Combining (3.4.2) for $\mu_{0}=\Phi(\nu)$ with (3.47), we have
$\mathbb{W}_{2}(\Phi(\mu),\Phi(\nu))\leq\sqrt{\tilde{c}_{2}}\text{\rm{e}}^{-\lambda t}\mathbb{W}_{2}(\Phi(\mu),\Phi(\nu))+\sqrt{\text{\rm{e}}^{(2+2K_{M})t}t}K_{I}\mathbb{W}_{2}(\mu,\nu),\mu,\nu\in\hat{\mathscr{P}}_{2,N}(\mathbb{R}^{d}).$
Let $\eta_{0}$ be given as in Theorem 3.3 and set
$\eta_{1}=\left(\inf_{t>\frac{\log{\sqrt{\tilde{c}_{2}}}}{\lambda}}\frac{\sqrt{\text{\rm{e}}^{(2+2K_{M})t}t}}{1-\sqrt{\tilde{c}_{2}}\text{\rm{e}}^{-\lambda t}}\right)^{-1}.$
So, when $K_{I}<\min(\eta_{0},\eta_{1})$, we can derive from the Banach fixed theorem that (3.44) has a unique invariant probability measure $\bar{\mu}\in\hat{\mathscr{P}}_{2}(\mathbb{R}^{2d})$.
Finally, we prove the exponential ergodicity in $\mathbb{W}_{2}$ and relative entropy. Similar to (3.47), we get
$\mathbb{W}_{2}(P_{t}^{\ast}\mu_{0},(\tilde{P}_{t}^{\bar{\mu}})^{\ast}\mu_{0})^{2}\leq\text{\rm{e}}^{(2+2K_{M})t}\int_{0}^{t}K_{I}^{2}\mathbb{W}_{2}(P_{s}^{\ast}\mu_{0},\bar{\mu})^{2}\text{\rm{d}}s,\ \ \mu_{0}\in\mathscr{P}_{2}(\mathbb{R}^{2d}).$
This together with (3.4.2) for $\mu=\bar{\mu}$ and Gronwall’s inequality yield
$\mathbb{W}_{2}(P_{t}^{\ast}\mu_{0},\bar{\mu})^{2}$
$\leq\left(\int_{0}^{t}2\tilde{c}_{2}\text{\rm{e}}^{-2\lambda s}2\text{\rm{e}}^{(2+2K_{M})s}K_{I}^{2}\text{\rm{e}}^{2\text{\rm{e}}^{(2+2K_{M})t}K_{I}^{2}(t-s)}\text{\rm{d}}s+2\tilde{c}_{2}\text{\rm{e}}^{-2\lambda t}\right)\mathbb{W}_{2}(\mu_{0},\bar{\mu})^{2}.$
So, by the same argument to derive (2.36), when $K_{I}$ is small enough, we can choose $\hat{t}>0$ and $\alpha\in(0,1)$ such that
$\mathbb{W}_{2}(P_{\hat{t}}^{\ast}\mu_{0},\bar{\mu})\leq\alpha\mathbb{W}_{2}(\mu_{0},\bar{\mu}),\ \ \mu_{0}\in\mathscr{P}_{2}(\mathbb{R}^{2d}).$
By the semigroup property $P_{t}^{\ast}P_{s}^{\ast}=P_{s+t}^{\ast}$, we get (3.41). Finally, by (3.4.1), (3.41) and the Talagrand inequality (3.34) with $\Phi(\mu)=\bar{\mu}$, the proof is completed by the same argument as (3.37). ∎